### v2

# Generating Verifiable Chain of Thoughts from Exection-Traces

Abstract.
Teaching language models to reason about code execution remains a fundamental challenge. While Chain-of-Thought (CoT) prompting has shown promise, current synthetic training data suffers from a critical weakness: the reasoning steps are often plausible-sounding explanations generated by teacher models, not verifiable accounts of what the code actually does. This creates a troubling failure mode where models learn to mimic superficially convincing but logically flawed reasoning patterns.
We address this by grounding CoT generation directly in program execution traces. Our pipeline instruments code to capture its dynamic behavior, then narrates these execution traces into natural language and factually-grounded rationales that are verifiable by design. This execution-grounded approach ensures every reasoning step reflects what the program computes, eliminating logical hallucinations at the source. We evaluate our method on code reasoning tasks, code generation and explanation tasks from HumanEval. Models trained on our bi-directional trace-grounded data achieve substantial improvements on reasoning tasks, with gains of up to 30 points on output prediction and 28 points on input prediction over base models, alongside competitive explanation and code generation performance. https://github.com/shailja-thakur/Verified-Code-CoT
Chain-of-Thought, Code Reasoning, Execution-Traces, Synthetic Data Generation, Program Semantics, Large Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering General programming languagesccs: Computing methodologies Natural language processing
1. Introduction
Recent advances in large language models have enabled code assistants that can generate, explain, and debug programs through natural language interaction. These systems—exemplified by GitHub Copilot, Claude Code, and similar tools—have demonstrated remarkable syntactic fluency. However, they struggle with tasks requiring deep reasoning about program behavior: debugging requires understanding how state evolves, program comprehension demands tracing execution paths, and code refactoring needs verifying semantic preservation. These reasoning tasks are fundamental to software engineering practice, yet current models often produce plausible but incorrect explanations of program behavior.
From a programming languages perspective, this reveals a deeper challenge: while LLMs are powerful static models of source code, they fundamentally lack a connection to the dynamic behavior of programs—how state evolves during execution. This gap limits their utility to that of sophisticated auto-completes rather than true reasoning partners for tasks like debugging and program analysis, which require a faithful understanding of a program’s operational semantics.
To bridge this gap, the research community has embraced fine-tuning models on Chain-of-Thought (CoT) data, which provides explicit, step-by-step rationales to guide the model’s reasoning process. However, the methods used to generate this data for code are often unsound. A powerful teacher model generates the rationale, but these reasoning steps are not tethered to a program’s actual execution trace. The result is a dataset of logical hallucinations—plausible but factually incorrect accounts of the program’s logic.
This challenge is especially acute in code reasoning, where small logical errors cascade into flawed understanding. Consider a model trained on a rationale claiming ”variable count is incremented in the loop”—when in fact, count is decremented. Such errors propagate during training, teaching the model to follow plausible but incorrect logical patterns. This lack of guaranteed soundness is a critical barrier for high-stakes applications. For debugging, a rationale that misrepresents variable state is misleading. For iterative self-refinement, feeding a model its own flawed reasoning causes compounding errors. For agentic workflows, a plan based on unsound understanding of a previous step can derail an entire task. Figure 1 illustrates this critical problem concretely.
Figure 1. Comparison of hallucinated vs trace-grounded CoT. Our approach translates pysnooper (pysnooper) execution trace (showing line numbers, variable states, state transitions) into natural language reasoning. Color mapping shows trace sections: yellow=init, blue=iterations, green=return. LLM hallucinates wrong comparison (arr[2] vs arr[1] instead of arr[3]) and wrong search direction.
Existing work uses execution in two ways. Benchmarks like HumanEval (chen2021evaluating) and methods like SemCoder (ding2024semcodertrainingcodelanguage) validate that the final program output is correct, but do not verify the logical integrity of intermediate reasoning steps. Formal verification methods can provide mathematical guarantees but are computationally prohibitive at scale. A critical gap exists: there is no practical method for validating each step in the reasoning chain—what we term rationale-step verification—that scales to large-scale training data generation. Additionally, existing pipelines lack fine-grained control over problem complexity and diversity.
Furthermore, effective code reasoning requires both forward inference (predicting outputs from inputs) and backward reasoning (understanding how outputs arose from inputs) (li2025codeio). This bi-directional capability is essential for debugging, code comprehension, and program analysis. However, current synthetic data generation methods do not systematically address this requirement, leading to models with asymmetric reasoning capabilities.
In this work, we bridge these gaps by introducing a new paradigm for data synthesis that achieves rationale-step verification at scale. We have developed a complete synthesis pipeline that generates CoT rationales by directly translating program execution traces into natural language. Specifically, we instrument a program to capture its dynamic execution state—recording variable values after each statement, control flow decisions, and state transitions derived from program semantic operations. This trace serves as an ground truth from which rationales are derived. Each step in the resulting reasoning chain can be verified against the trace to confirm: (1) variable values match recorded runtime values, (2) state transitions correctly reflect how variables changed between statements, and (3) control flow accurately describes which branches executed. This ensures rationales are verifiable by design, eliminating the logical hallucinations that plague other methods.
Figure 2. An overview of our three-stage data synthesis pipeline. Stage A (Concept Sourcing & Synthesis): Generates candidate concepts, code, and tests from raw documents. Stage B (Verification & Clustering): Uses our execution-based Dual Agreement algorithm to identify and rank the highest-quality solution-test pairs. Stage C (CoT Generation): Uses the verified artifacts and their execution traces to produce the final, bi-directional conversational data, complete with trace-grounded rationales and feedback.
Table 1. Comparison of features in related works on code reasoning data and methodologies. Our work is the first to systematically combine execution-grounded verification with bi-directional CoT generation at scale.
Work / Method Execution-Trace Natural Language Bi-Directional Verifies Scalable Data
Grounded CoT CoT CoT Output Synthesis
REVTHINK (chen2024reverse)
TRACED (ding2024traced)
CodeI/O (li2025codeio)
rStar-Coder (liu2025rstarcoderscalingcompetitivecode)
Jung et al. (jung2025codeexecutiongroundedsupervision)
SemCoder (ding2024semcodertrainingcodelanguage)
Our Work
Furthermore, our concept-first pipeline provides fine-grained control over data quality. Rather than relying on existing code corpora, it synthesizes problems from abstract programming concepts (e.g., ”pass-by-reference semantics,” ”dynamic programming”) extracted from technical literature. This allows us to systematically control problem difficulty and ensure diverse coverage of algorithmic and mathematical domains. Combined with bi-directional trace generation, this provides a complete foundation for training models on both forward reasoning (predicting outputs) and backward reasoning (diagnosing how outputs arose), with all rationales grounded in verifiable execution.
Our contributions are fivefold:
  1. (1)
    Chain of Thought Step Verification at Scale: We introduce a scalable execution-grounded synthesis pipeline that achieves rationale-step verification.
  2. (2)
    Concept-First Curriculum Synthesis: A curriculum-driven synthesis that provides control over problem complexity and diversity, enabling systematic coverage from basic constructs to advanced algorithms.
  3. (3)
    Comprehensive Empirical Validation: Through systematic validations, we demonstrate that execution-groundedness and bi-directional nature of CoTs directly determine model performance.
  4. (4)
    Open-Source Pipeline: Complete synthesis infrastructure including concept extraction, hierarchical generation, Dual Agreement verification, and trace-to-rationale translation to facilitate reproducible research in verified reasoning for code.
The remainder of this paper is organized as follows. We first review related work in code reasoning and synthetic data generation. Next, we detail our three-stage data synthesis pipeline, followed by our experimental design and a thorough analysis of the results. We conclude with a discussion of our findings and future research directions.
2. Related Work
Our work builds upon several interconnected research areas in code reasoning, execution-based program understanding, and synthetic data generation. Chain-of-Thought (CoT) prompting (wei2022chain) has proven effective for improving reasoning, but its extension to code faces unique verification challenges. While methods like Program-of-Thought (PoT) (chen2022program) generate executable snippets, they lack systematic verification of the reasoning process itself. Building on this foundation, CodeI/O (li2025codeio) takes a significant step forward by transforming code patterns into natural language CoT rationales through input/output prediction tasks. By training models to predict inputs and outputs given code, CodeI/O exposes models to universal reasoning primitives while decoupling reasoning from syntax. However, CodeI/O focuses on prediction tasks rather than generating complete reasoning datasets from execution traces.
Verification Approaches in Code Generation. The term ”verification” in code generation literature often refers to different concepts. We identify three distinct approaches:
  1. (1)
    Outcome Verification validates only the final program output against expected results. This is the most common approach, used by benchmarks like HumanEval (chen2021evaluating) and methods like SemCoder (ding2024semcodertrainingcodelanguage) and rStar-Coder (liu2025rstarcoderscalingcompetitivecode). While effective for validating solutions, it offers no guarantee about the logical integrity of intermediate reasoning steps.
  2. (2)
    Formal Verification uses mathematical proofs (e.g., symbolic execution, SMT solvers) to guarantee program correctness. While sound, this approach is computationally heavyweight and impractical for large-scale training data generation.
  3. (3)
    Rationale-Step Verification (our contribution) validates each reasoning step in the CoT chain. By generating rationales directly from a program’s execution trace—a factual record capturing variable values, state transitions, and control flow—we verify that each reasoning step correctly describes runtime states and semantic operations. This verifies the logical integrity of the entire reasoning process, not just the outcome.
Recent work like TRACED (ding2024traced) and Execution Tuning (E.T.) (armengol2025execution) has explored using execution traces for pre-training, but these methods do not generate natural language rationales with step-by-step verification. Our work is the first to systematically achieve rationale-step verification at scale by translating execution traces into verifiable natural language CoT. Conceptually closest to our trace translation, Jung et al. generate natural language CoT from execution traces but focus on single-direction reasoning rather than a systematic bi-directional pipeline.
Comparison to SemCoder and rStar-Coder. SemCoder and rStar-Coder also generate synthetic CoT data, but our methodology differs in two aspects. First, while SemCoder generates rationales via an LLM’s explanation and verifies only output, we perform a direct translation of execution traces, eliminating logical hallucinations. Second, unlike SemCoder’s code-first approach, our concept-first approach provides fine-grained control over complexity and diversity by synthesizing problems from abstract technical concepts. Similarly, rStar-Coder excels at verifiably correct solutions, while we focus on creating verifiably correct reasoning processes.
Bi-Directional Reasoning. The principle of bi-directional reasoning has been explored in code by CodeI/O (li2025codeio) through I/O prediction tasks, and in other domains by FOBAR (jiang2024forward) and REVTHINK (chen2024reverse). However, these approaches focus on prediction tasks rather than generating complete trace-grounded rationales. Other verification techniques like Self-Verification (weng2023large) and RCoT (xue2023rcot) rely on complex checking procedures, while our deterministic trace-based method provides guaranteed correctness.
As summarized in Table 1, no existing work provides a systematic pipeline for generating bi-directional CoT datasets where every reasoning step is grounded in verifiable execution. Our work addresses this gap, providing a foundation for training robust code reasoning capabilities.
3. Data Synthesis Pipeline
Our data synthesis pipeline (Figure 2) is a multi-stage process designed to generate high-fidelity, bi-directional CoT data where every reasoning step is anchored in ground-truth program execution.
3.1. Stage A: Concept Sourcing and Curriculum-Driven Synthesis
Rather than generating code from simple prompts, our pipeline begins by building a curriculum of programming concepts derived from high-quality sources. This curriculum-driven approach ensures the resulting problems are complex, diverse, and grounded in established knowledge.
3.1.1. Document Processing and Concept Extraction
Our pipeline processes a diverse corpus of permissively-licensed technical literature, including books from the StarCoder2-documentation dataset (starcoder2-doc) and curated programming resources spanning basic to advanced topics. Rather than relying on raw PDF text extraction, which produces noisy output, we employ Docling (Docling), a document understanding framework that renders PDFs into clean, structured markdown. This preprocessing preserves semantic structure (headings, code blocks, lists), removes pagination artifacts, and maintains proper formatting. The cleaned text is chunked into 4000-character segments with sliding overlap to prevent concept fragmentation.
3.1.2. Hybrid Concept Identification
For each text chunk, we employ a three-stage hybrid extraction strategy balancing recall with precision:
Stage 1: Statistical Keyword Extraction. We use spaCy with PyTextRank, a graph-based ranking algorithm, to identify candidate concepts. PyTextRank applies the TextRank algorithm (adapted from PageRank) to the chunk’s lemma graph, where nodes represent terms and edges represent co-occurrence relationships. This produces a ranked list of phrases based on their centrality. However, this captures non-technical terms (e.g., page numbers, dates) and misses implicit concepts discussed narratively.
Stage 2: LLM-Based Filtering and Augmentation. To address these limitations, we prompt an LLM teacher model (Qwen2.5-Coder-7B-Instruct) with a filtering prompt that: (1) removes book metadata and non-Python-specific terms, (2) completes incomplete phrases, and (3) identifies implicit concepts present in text but not captured statistically. This hybrid approach yields approximately 15,000 initial concepts.
Why This Hybrid Approach? Pure NLP methods lack semantic understanding and generate noise; pure LLM extraction is expensive and may hallucinate concepts not in text. Our hybrid leverages the efficiency of statistical methods for recall and LLM semantic reasoning for precision.
3.1.3. Rigorous Concept Deduplication and Quality Scoring
The initial 15,000 concepts contain substantial redundancy and quality variance. We apply multi-stage quality control:
Deduplication. We normalize concepts (lowercasing, lemmatizing) and perform string similarity matching to cluster near-duplicates, selecting the most complete variant as canonical.
LLM-Based Quality Scoring. For each deduplicated concept, we prompt the teacher model to score along two dimensions: (1) Difficulty (1-5): algorithmic complexity required for implementation, and (2) Relevance (1-5): ability to inspire diverse, non-trivial problems. We retain only concepts scoring $\geq 3$ on both dimensions, yielding approximately 8,000 high-quality seed concepts.
Why Difficulty Filtering Matters. Preliminary experiments revealed that naively sampling produces a distribution heavily skewed toward trivial problems. This is problematic because: (1) trivial problems do not challenge multi-step reasoning, (2) resulting CoT rationales are too short for meaningful supervision, and (3) downstream verification struggles with trivial problems where nearly all candidates pass all tests. Difficulty filtering biases the curriculum toward problems requiring algorithmic thinking.
[…]
¡Instruction¿
Given a weighted graph, find the most cost-effective path between two nodes.
¡Function¿
def find_cheapest_path(graph, start, end): ...
¡Question¿
For the given network, what is the minimum cost and route from node ’A’ to ’F’?
¡Instruction¿
Given a weighted graph, find a starting node for a cost-effective path.
¡Function¿
def find_cheapest_path(graph, start, end): ...
¡Question¿
What start node could result in a path to ’F’ with total cost of 7?
¡CoT¿ (1) Initialize distances: ’A’ to 0, all others to $\infty$. Priority queue starts with (’A’, 0). (2) Extract ’A’ from queue. Update distances for neighbors ’B’ and ’C’; add to queue. (3) Continue extracting minimum, updating neighbors until destination ’F’ reached. (4) Backtrack from predecessors to reconstruct path. Final cost is 7. ¡Predicted Output¿
([’A’, ’C’, ’E’, ’F’], 7)
¡CoT¿ (1) The final path to ’F’ has cost 7. To deduce start, trace predecessors backwards. (2) The predecessor of ’F’ in cost-7 path must be ’E’. Edge cost from ’E’ to ’F’ is 2, so cost to reach ’E’ is 5. (3) Continue backtracking through ’C’, which confirms starting at ’A’ achieves total cost of 7. ¡Predicted Input¿
’A’
Figure 3. Bi-directional CoT data format examples. Blue tags denote prompt components (instruction, function, question); green regions show model’s trace-grounded reasoning and prediction. Code appears in monospace, reasoning in normal text.
Function Signature   — Class Signature
solution(freq_list: list[tuple[str,int]]) -> encoding: dict[str, str] Extracted Metadata:
• Function: solution
• Param name: freq_list
• Param type: list[tuple[str,int]]
• Return name: encoding
• Return type: dict[str,str]
Class HuffmanTree: __init__(self, freq_list: list[tuple[str,int]]) -> None build_tree(self) -> tree: tuple get_encoding(self) -> codes: dict[str,str] Extracted Metadata:
• Class: HuffmanTree
• Constructor param: freq_list: list[tuple[str,int]]
• Method: build_tree, returns tree:tuple
• Method: get_encoding, returns codes:dict[str,str]
Figure 4. Signature format templates with extracted metadata. Generated signatures specify function/class names, parameter names and types, and return variable names with types (e.g., tree: tuple, codes: dict). This metadata is stored to validate code/test consistency and prevent hallucinated names or type mismatches.
Correct Format   — Prohibited Format
def test_basic():
  # Test basic case
  assert
solution([1,2,3], 2) == [1]
def test_wrong():
  # Variable outside assert
  lst = [1, 2, 3]
  assert
solution(lst, 2) == [1]
Figure 5. Test format requirements. Correct format (left) enables clean I/O extraction for trace generation with direct function calls in assert statements, while prohibited format (right) complicates trace analysis with intermediate variable assignments.
3.1.4. Hierarchical Problem Synthesis
For each high-quality concept, we synthesize a complete problem artifact set through a carefully ordered five-step prompting pipeline. The ordering is designed to constrain the solution space at each step, preventing type and naming inconsistencies.
Step 1: Instruction Generation. Given a concept and its description, we prompt the teacher model to generate distinct natural language problem instructions. The prompt is engineered for diversity along three axes: (1) problem domain (mathematics, finance, data processing, algorithms), (2) computational approach (iterative, recursive, dynamic programming, functional), and (3) difficulty (”medium” targeting 30-60 LOC, ”hard” targeting 50-100+ LOC). The prompt includes explicit anti-overlap guidance to avoid paraphrased versions of the same problem.
Step 2: Signature Generation. Before generating any code or tests, we prompt the LLM to analyze the instruction and produce a formal signature skeleton specifying: (1) implementation type (standalone function vs. class), (2) function signature with parameter types and return type, or (3) class signature with constructor and methods. The signature prompt uses a strict formatting template enforced via few-shot examples, as shown in Figure 4. From each signature, we extract and store metadata including function names, parameter names and types, and crucially, return variable names with their types (e.g., encoding: dict[str, str]). This metadata serves as ground truth for validating consistency across code, tests, and generated CoT rationales. This signature-first design is deliberate: once input/output types and function names are fixed, subsequent code generation cannot hallucinate different types or names, eliminating import failures and ensuring test-code alignment.
Step 3: Code Generation. With the signature locked, we generate candidate solutions per instruction by prompting with both the instruction and signature. The prompt explicitly instructs: (1) strictly adhere to the signature, (2) vary computational approach across implementations, and (3) ensure completeness.
Step 4: Test Scenario Identification. We prompt the LLM to analyze the problem and produce a concise list of test scenarios (e.g., ”Test basic functionality”, ”Test empty input”). This improves test coverage by encouraging reasoning about requirements before writing code.
Step 5: Test Generation. We generate unit tests (3 test suites per instruction, each with up to 10 test functions) using a meticulously engineered prompt enforcing strict structural constraints. Each test is a top-level function (not a class method) containing exactly one assert statement with direct function calls using inline arguments—no variable assignments. Figure 5 illustrates the correct format versus prohibited patterns. These constraints are critical for Stage B granular pass/fail clustering and Stage C clean trace extraction, while enabling parallel sandboxed execution.
3.2. Stage B: Execution-Based Verification and Agreement Clustering
The hierarchical generation in Stage A, while carefully constrained, produces noise: some candidate solutions are incorrect (due to LLM generation errors), some tests are malformed or have wrong expected outputs, and some instruction-signature pairs are ambiguous. Before generating expensive CoT rationales in Stage C, we must identify and discard low-quality artifacts. We adapt the Dual Agreement verification methodology from CodeT (chen2022codetcodegenerationgenerated), originally designed for single-solution selection, into a scalable batch filtering algorithm.
3.2.1. Mass Execution and Pass/Fail Matrix Construction
For each problem (identified by task_id), we have $m$ candidate solutions and $n$ candidate tests. We execute all $m\times n$ solution-test pairs in a secure, sandboxed environment (Docker containers with resource limits) to construct a binary pass/fail matrix $M\in\{0,1\}^{m\times n}$, where $M[i,j]=1$ if solution $i$ passes test $j$, and 0 otherwise. Execution failures (timeouts, runtime errors, assertion failures) are all treated as $M[i,j]=0$. This brute-force execution is computationally expensive ($O(m\cdot n)$ per task) but embarrassingly parallel.
3.2.2. Dual Agreement Clustering
The core insight of Dual Agreement is: If a large set of independently generated solutions all pass the same large set of independently generated tests, it is statistically unlikely that both the solutions and tests are incorrect. We formalize this as follows:
Clustering by Test Agreement. We partition the $m$ candidate solutions into clusters $\{C_{1},C_{2},\ldots,C_{k}\}$ such that all solutions within a cluster $C_{i}$ have identical pass/fail patterns across all $n$ tests. Formally, solutions $s_{a}$ and $s_{b}$ are in the same cluster iff:
$\forall j\in[1,n]:M[a,j]=M[b,j]$
This is implemented via hash-based grouping: we compute a fingerprint for each solution’s test pass pattern (bit vector hash), then group solutions by fingerprint.
Cluster Scoring. For each cluster $C_{i}$, let $T_{p}(C_{i})$ denote the set of tests that all solutions in $C_{i}$ pass. We assign a quality score:
$\text{Score}(C_{i})=|C_{i}|\times|T_{p}(C_{i})|$
This score captures two dimensions: (1) Solution agreement ($|C_{i}|$): If many independently generated solutions have identical behavior, they likely implement the same correct algorithm. (2) Test coverage ($|T_{p}(C_{i})|$): If they pass many tests, the tests likely have correct expected outputs.
Best Cluster Selection. For each task, we rank clusters by score and select the single highest-scoring cluster. We extract one canonical solution from this cluster (selecting the pair with maximum code coverage) and its associated passing tests. Algorithm 1 formalizes this procedure. This method effectively filters multiple noise sources: incorrect solutions land in low-scoring clusters due to failing many tests (low $|T_{p}|$), malformed tests reduce scores but cannot eliminate correct clusters if other valid tests exist, and ambiguous instructions cause solutions to fragment into multiple small clusters (low $|C_{i}|$), with the highest-scoring cluster representing the most common interpretation.
Algorithm 1 Dual Agreement Clustering
0: Solutions $S=\{s_{1},\ldots,s_{m}\}$, Tests $T=\{t_{1},\ldots,t_{n}\}$
0: Verified solution $s^{*}$ and test suite $T^{*}$
1: Execute all $m\times n$ pairs to build matrix $M\in\{0,1\}^{m\times n}$
[…]
3:for each solution $s_{i}\in S$ do
4:  $\textit{pattern}_{i}\leftarrow M[i,:]$ // Test pass/fail pattern
5:  Add $s_{i}$ to cluster with matching pattern in clusters
6:end for
7:for each cluster $C_{i}\in\textit{clusters}$ do
8:  $T_{p}(C_{i})\leftarrow\{t_{j}:\forall s\in C_{i},M[s,j]=1\}$ // Commonly passed tests
9:  $\textit{Score}(C_{i})\leftarrow|C_{i}|\times|T_{p}(C_{i})|$
10:end for
[…]
14:return $(s^{*},T^{*})$
3.2.3. Theoretical Justification: Why Multiple Candidates Matter
The effectiveness of Dual Agreement relies critically on generating multiple diverse candidates. While a single solution-test pair provides no signal about correctness, multiple independently generated candidates enable statistical consensus detection through two complementary properties.
Unlikely Collision Property. The probability that $K$ independently incorrect solutions would all pass the same incorrect test suite decreases exponentially with $K$. Formally, if each incorrect solution has probability $p$ of coincidentally passing an incorrect test, then the probability that all $K$ incorrect solutions pass is $p^{K}$. For reasonable values (e.g., $p=0.3$, $K=5$), this yields $p^{K}=0.00243$, making such false consensus highly unlikely. In contrast, $K$ correct solutions deterministically pass correct tests. This means large $|C_{i}|$ (solution agreement) and large $|T_{p}(C_{i})|$ (test coverage) together provide strong correctness signals—clusters with both high $|C_{i}|$ and high $|T_{p}|$ survive the scoring threshold, filtering out isolated or weakly-validated artifacts.
We can formalize why this scoring creates clear separation between correct and incorrect clusters. Let $C_{correct}$ be the cluster of correct solutions and $C_{error}$ be any cluster of incorrect solutions. For the correct cluster:
$\mathbb{E}[\text{Score}(C_{correct})]=\mathbb{E}[|C_{correct}|]\times n_{valid}$
where $n_{valid}$ is the number of valid tests. For an incorrect cluster:
$\mathbb{E}[\text{Score}(C_{error})]\leq\mathbb{E}[|C_{error}|]\times n_{valid}\times(1-\delta)$
where $\delta>0$ represents the fraction of tests that expose the bug. When we have $K$ diverse candidate solutions with $m$ correct among them, the expected score ratio becomes:
$\frac{\mathbb{E}[\text{Score}(C_{correct})]}{\mathbb{E}[\text{Score}(C_{error})]}\geq\frac{m}{(K-m)(1-\delta)}$
As $K$ increases with sufficient diversity, $m$ (correct solutions) grows while incorrect solutions fragment into multiple small clusters with different error patterns, each with low $|C_{error}|$. This multiplicative scoring creates exponential separation, allowing correct clusters to dominate even with noisy candidates. However, this requires diversity: generating varied candidates via temperature sampling, prompt variation, and varied computational approaches in Stage A ensures that correct implementations converge while incorrect ones fragment. The effectiveness of this approach is empirically validated in Section 5.1.
3.3. Stage C: Execution-Grounded CoT Generation
The final stage transforms each verified solution-test pair into a rich, conversational training dataset where every reasoning step is grounded in the program’s actual execution. This is not a single-prompt process, but a meticulous five-step automated pipeline.
3.3.1. Step 1: Execution Trace Generation
For each verified solution and selected passing test cases, we generate a ground-truth execution trace capturing the program’s dynamic behavior. We use pysnooper(pysnooper), a lightweight Python tracing library, to instrument the target function (identified via signature metadata from Stage A).
Instrumentation Strategy. Rather than tracing the entire program (which includes test framework overhead), we selectively instrument only the function under test. This approach is robust against issues like recursion or complex interprocedural calls because the structural constraints imposed during Code Generation (Stage A)—which enforce a flattened logic without nested functions—guarantee that all core computation resides within the instrumented scope. For a function solution, we decorate it with @pysnooper.snoop(output=trace_file), then execute the test case in a sandboxed subprocess with a strict timeout (10 seconds). This produces a line-by-line trace file detailing: (1) Call events: Function entry with input arguments, (2) Line execution: Each line number executed in sequence, (3) Variable modifications: Every assignment with variable name, old value, and new value (capturing state transitions), and (4) Return events: Function exit with return value.
This trace is the immutable ground truth: a factual record of the program’s operational semantics—what values variables held at each point, how those values changed from statement to statement, and which execution paths were taken. It is not an LLM’s explanation or interpretation, but the actual runtime behavior.
3.3.2. Step 2: Trace Sanitization and Normalization
Raw pysnooper logs contain formatting artifacts unsuitable for model consumption: ANSI color codes, timestamp prefixes, file path prefixes, and redundant whitespace. A dedicated sanitization script strips these artifacts via regex-based cleaning, producing a clean, plain-text trace representation. This is critical because LLMs are sensitive to formatting: spurious tokens can degrade generation quality.
3.3.3. Step 3: Ground-Truth I/O Extraction and Question Generation
Before generating CoT rationales, we establish the precise input and output for each test case. For simple test functions like def test_basic(): assert solution([1, 2], 2) == [1], we use regex parsing to extract input = [1, 2], 2 and expected_output = [1]. For complex tests involving setup logic, we prompt an LLM to analyze the test code and extract the effective input/output pair. With ground-truth I/O established, we then prompt an LLM to generate a pair of natural language questions: (1) Forward question: ”Given the input [1, 2, 3] and target=2, what does the function return?”, and (2) Backward question: ”What input would cause the function to return [1]?” The prompt explicitly instructs natural phrasing to add linguistic diversity to the dataset.
3.3.4. Step 4: Trace-Grounded Chain-of-Thought Generation
This is the core of Stage C, where we use the sanitized trace as the factual basis for generating reasoning chains.
Forward CoT Generation: Narrating the Trace. We prompt an LLM with: (1) the instruction and function code, (2) the forward question, and (3) the sanitized execution trace. The prompt instructs: ”Narrate the execution trace step-by-step, explaining how the input is transformed into the output. Extract the <Predicted Output> you deduce from the trace.” Critically, the model is not generating an explanation from scratch; it is translating a factual trace into natural language. Each statement in the rationale can be verified against the trace: variable values mentioned must match the recorded runtime values, state transitions described must correspond to logged changes between statements, and the sequence of operations must follow the actual control flow path executed.
Backward CoT Generation: Deductive Reasoning from the Trace. For the backward direction, we prompt the model with the same trace but ask it to perform deductive reasoning: ”Given the final output state in the trace, explain how this output could only have been reached from the initial input. Extract the <Predicted Input> you derive.” This forces the model to reason backwards through the trace, explaining how the final state constrains the initial input.
Why This Trace-Grounding Eliminates Hallucination. Traditional self-instruct methods prompt an LLM to explain code without execution, leading to ”plausible hallucinations” where the explanation sounds correct but contains subtle factual errors (e.g., claiming a variable was incremented when it was decremented). By anchoring the rationale in a trace, we guarantee factual correctness: (1) Variable values are verifiable: Each value mentioned in the rationale is recorded at runtime and can be cross-referenced against the trace log. (2) State transitions are verifiable: Claims about how variables changed between statements (e.g., ”the counter increases from 0 to 1”) can be confirmed by comparing consecutive trace entries showing the variable’s old and new values. (3) Control flow is verifiable: Descriptions of which branches or loops executed can be validated against the trace’s record of which lines were executed and in what order. This ensures the entire causal chain—from initial state through semantic operations to final state—is factually grounded.
3.3.5. Step 5: Final Assembly into Conversational Format
The generated components (questions, CoTs, predicted I/O) are assembled into structured training examples. We create three versions of the dataset: (1) Forward-only: Each sample contains <Instruction>, <Function>, <Forward Question>, <CoT>, <Predicted Output>; (2) Backward-only: Each sample contains <Instruction>, <Function>, <Backward Question>, <CoT>, <Predicted Input>; (3) Bi-directional: Each sample contains both forward and backward turns in a single multi-turn conversation, teaching the model to reason in both directions.
4. Experimental Design
Our experiments are designed to rigorously quantify the benefits of our execution-grounded data and to determine the optimal configurations for fine-tuning.
4.1. Models and Benchmarks
Our model selection was deliberately designed to test our method across two open-source foundation models of a similar scale but with distinct pre-training philosophies. We selected granite-3.3-8b-base (granite3.3-base), an enterprise-grade architecture trained exclusively on permissively licensed data, which validates our approach for real-world applications where data provenance is critical. Alongside it, we chose Qwen2.5-Coder-7B (hui2024qwen2), a state-of-the-art, code-native specialist, to test if our data provides a reasoning signal capable of enhancing even expert models. The ability to improve both a safe generalist and a top specialist demonstrates the fundamental and broad applicability of our method. We evaluate all models on LiveCodeBench (Execution) (jain2024livecodebench) and CruxEval (Input/Output) (gu2024cruxeval), benchmarks specifically chosen for their focus on code execution reasoning. We finetune the base models for 10 epochs with a starting LR of 2e-6. More details about the training setup are included in the Appendix.
4.2. Data Curation and Preparation
The foundation of our experiments is a large-scale, high-quality dataset generated by our synthesis pipeline.
  1. (1)
    Concept Sourcing and Scaling: The pipeline began by ingesting a diverse set of permissively-licensed documents, including those from the ”Free Programming Books” source in the StarCoder2-documentation dataset (starcoder2-doc) and curated lists of programming books. These sources cover a wide range of topics, from core Python language features and data structures to advanced algorithms. This process yielded an initial set of approximately 15,000 deduplicated concepts. From this seed, we hierarchically synthesized a large pool of programming problems: for each concept, we generated multiple distinct instructions, and for each instruction, multiple candidate solutions and unit tests.
    4.2.1. Hyperparameter Selection for Data Generation
    To determine the optimal number of candidate solutions and unit tests per problem, we conducted systematic experiments with varying configurations. We evaluated different numbers of candidate solutions (5, 10, 15, 20) and unit test cases (5, 10, 20, 30, 40, 50) per instruction. Based on our ablation studies balancing quality and computational cost, we selected 5 candidate solutions and 30 unit tests per problem. The 30 unit tests were generated in 3 sets of 10 unit tests each, as described in Section 3.1, ensuring comprehensive coverage of edge cases and core functionality while maintaining diversity through multiple independent generation rounds.
  2. (2)
    Verification and Deduplication: This large, raw set of candidates was passed through our execution-based verification stage (Stage B), resulting in approximately 85,000 code-test pairs that were confirmed to be functionally correct. A final deduplication pass on similar code snippets and test logic yielded our initial master dataset.
    CoT Formatting: Each sample in the master set was formatted into three distinct reasoning structures. A forward CoT sample presents a question like Given input X, what is the output? followed by a trace-narrated rationale. A backward CoT sample asks "What input could produce output Y?" followed by a deductive, trace-grounded rationale. Figure 3 illustrates both formats showing the complete structure: instruction, function, question, step-by-step trace-grounded reasoning, and predicted output/input. Finally, a bi-directional sample combines these into a single, multi-turn conversational format.
  3. (3)
    Difficulty-Based Subsetting: From the master dataset of 85k code-test pairs, we created three distinct subsets for our ablation studies, based on different filtering strategies:
    • Full Set (54k): This is the complete, verified, and deduplicated dataset, representing the broadest collection of concepts.
    • Model Answerability Set (25k): This high-quality subset was created via a novel filtering strategy: we kept only the problems from the 54k deduplicated set that an LLM failed to solve correctly. This creates a targeted curriculum focused on the model’s specific weaknesses.
    • Content-Rated Difficulty Set (18k): This set was further refined from the 25k set by using an LLM to rate the conceptual difficulty of each problem, retaining only those rated as ”medium” or ”hard”.
This curation process provides two key dimensions for our experiments: the data curation strategy (54k vs. 25k vs. 18k) and the reasoning direction (forward, backward, or bi-directional).
4.3. Ablation Study Plan
Our experiments are designed as a sequential, top-down filter to efficiently identify the data configuration. While testing every combination of our data subsets and reasoning formats would be ideal, such a full factorial experiment is computationally infeasible. Our sequential approach is a practical and methodologically sound alternative that allows us to isolate the impact of our two primary contributions: difficulty-filtered subsets and CoT directionality. We operate on the reasonable assumption that the data subset demonstrating the highest quality for forward CoT will also be the most effective foundation for the other CoT formats. All experiments are conducted in a Supervised Fine-Tuning (SFT) setting.
  1. (1)
    Data Curation Ablation: We first aim to identify the optimal data subset. We fine-tune each base model on the 54k, 25k, and 18k datasets using only the forward CoT samples and compare their performance.
  2. (2)
    CoT Direction Ablation: Using the winning data subset from the previous stage, we then investigate the impact of the CoT format by fine-tuning models on the forward-only, backward-only, and the full bi-directional versions of that dataset.
  3. (3)
    Model Generalization: By performing this two-stage process on both Granite and Qwen-Coder, we validate the generalization of our findings across different model architectures.
5. Results and Analysis
We conduct a series of systematic ablation studies to evaluate the effectiveness of our data synthesis pipeline and determine the optimal training configurations. Our experimental design follows a sequential, top-down filtering approach: we first identify the best-performing data curation strategy, then use that dataset to determine the most effective reasoning direction. This entire process is performed on Granite-3.3-8b-base, and then repeated for Qwen2.5-Coder-7B to validate the generalizability of our findings. We evaluate all models on LiveCodeBench (Execution) and CruxEval (Output and Input prediction).
5.1. Validating Dual Agreement
Before presenting the main experimental results, we validate the effectiveness of our Dual Agreement verification approach (Stage B) through a controlled study on 5,000 programming problems with ground-truth correct solutions. Using the hyperparameter configurations described in Section 5.1.1, we measured both consensus scores and alignment with ground truth to quantify how the dual dimensions of solution agreement and test coverage affect verification reliability.
5.1.1. Consensus Score Scaling with Candidates
Figure 6 shows a heatmap of the percentage of problems achieving high consensus scores (Score $\tau$ threshold) as a function of the number of solutions (x-axis) and test cases (y-axis). The heatmap reveals a clear pattern: the bottom-right region (high solutions, high tests) shows consistently high consensus rates (70-90% of problems), while the upper-left region (few solutions, few tests) shows near-zero consensus rates. This validates our hypothesis that both multiple diverse solutions and comprehensive test suites are necessary for reliable verification. Notably, increasing solutions from 5 to 20 yields a 3x increase in problems with high consensus scores when test coverage is adequate ($\geq 20$ tests). This empirically confirms the theoretical exponential decay property: more candidates exponentially reduce the probability of false consensus.
Figure 6. Heatmap showing percentage of problems achieving high consensus scores as a function of number of candidate solutions (x-axis) and test cases (y-axis). The bottom-right region (many solutions, many tests) achieves 70-90% consensus, while the top-left (few solutions, few tests) shows near-zero consensus, validating that both dimensions are necessary for reliable verification.
Table 2. Comprehensive evaluation results across models, datasets, and training configurations on CruxEval and LiveCodeBench-Exec benchmarks. The experiments follow a top-down approach, with the winning configuration from one stage used in the next. Best results for each stage are highlighted in bold. And best result across the exp. stages are highlighted in grey.
Model CoT Direc. Data Subset LCB-Exec CruxEval Output CruxEval Input
Pass@1 Pass@1 Pass@5 Pass@1 Pass@5
Granite-3.3-8B-base Base (default) N/A 18.3 15.5 25.3 14.3 28.4
Granite-3.3-8B (FT) Fwd 18k 43.5(+25.2) 36.1(+20.6) 58.2(+32.9) 35.8(+21.5) 57.9(+29.5)
25k (best) 44.9(+26.6) 42.7(+27.2) 64.7(+39.4) 40.2(+25.9) 63.5(+35.1)
54k 34.1(+15.8) 28.9(+13.4) 55.2(+29.9) 28.8(+14.5) 54.9(+26.5)
[…]
Bwd 25k 35.4(+17.1) 39.3(+23.8) 61.3(+36.0) 41.5(+27.2) 64.8(+36.4)
[…]
Qwen2.5-Coder-7B Base (default) N/A 46.3 45.3 52.12 47.5 49
Qwen2.5-Coder-7B (FT) Fwd 18k 66.9(+20.6) 58.4(+13.1) 75.5(+23.4) 57.2(+9.7) 69.4(+20.4)
25k 67.0(+20.7) 57.5(+12.2) 73.9(+21.8) 59.8(+12.3) 67.5(+18.5)
54k (best) 66.5(+20.2) 58.6(+13.3) 76.0(+23.9) 60.5(+13.0) 68.3(+19.3)
Fwd 25k 67.0(+20.7) 57.5(+12.2) 73.9(+21.8) 59.8(+12.3) 67.5(+18.5)
Bwd 25k 57.5(+11.2) 50.4(+5.1) 69.8(+17.7) 61.2(+13.7) 69.1(+20.1)
[…]
Note: All Pass@k scores are reported as percentages. Green subscript values show improvement over the Base (Pre-trained) baseline. Abbreviations: Filt. = Filtered, perf. = performing, Fwd: Forward only CoT samples, Bwd: Backward only CoT samples, FT: Supervised fine-tuned.
5.2. Impact of Data Curation
Our first primary ablation investigates the impact of data quality and size by fine-tuning models on three forward-only data subsets: a complete 54k sample set, a higher-quality 25k set filtered for correctness, and a challenging 18k set filtered for difficulty.
For the Granite-3.3-8b model, the results are definitive. As shown in Table 2, the 25k correctness-filtered dataset substantially outperforms the others across all benchmarks. On LiveCodeBench, it achieves a score of 44.9%, an absolute gain of +26.6 over the base model, and significantly outperforms the larger 54k set (34.1%). This powerful improvement from a targeted, high-quality dataset indicates that for complex reasoning, data verifiability is far more impactful than sheer volume.
When repeating the experiment with Qwen2.5-Coder-7B, all three fine-tuned models dramatically outperform the base model. For instance, the 25k dataset boosts the LiveCodeBench score from 46.3% to 67.0% (+20.7). Unlike with Granite, the performance across the three data subsets is competitive. The 25k dataset achieves the highest score on LiveCodeBench, while the 54k dataset has a slight edge on the CruxEval benchmarks. Given the consistently strong performance of the 25k set, we selected it as the winning configuration for all subsequent experiments.
Figure 7. Performance comparison against SOTA baselines (ding2024semcodertrainingcodelanguage) Our model’s results are highlighted with a hatch pattern. Abbreviations: CL: CodeLlama, SC2: StarCoder2, DSCoder: DeepSeekCoder, MCoder: MagicCoder, -Py: Python, -Inst: Instruct.
5.3. Impact of Reasoning Direction
Using the winning 25k dataset from the previous stage, we evaluated the impact of our novel bi-directional data format. We compared models trained on forward-only, backward-only, and the complete bi-directional datasets.
For the Granite-3.3-8b model, the results in Table 2 highlight the benefits of bi-directional training. The model trained on the bi-directional dataset achieves the highest scores on both CruxEval Output (45.7% pass@1) and Input (42.1% pass@1). This represents a final performance gain of +30.2 on CruxEval Output over the base model, demonstrating that teaching cause-to-effect and effect-to-cause reasoning provides a synergistic improvement.
This finding is further confirmed with the Qwen2.5-Coder-7B model. The bi-directionally trained model once again emerges as the top performer, achieving the best results on LiveCodeBench (68.2%) and CruxEval Input (61.9% pass@1). For this configuration, final score on LiveCodeBench represents a +21.9 point gain over the base model. To contextualize this performance, Figure 7 compares our best model against a range of competitive foundation models. Our fine-tuned model establishes a leading score on LiveCodeBench-Execution and demonstrates highly competitive performance across all reasoning benchmarks.
5.4. Does the method used to generate and verify chain-of-thought rationales fundamentally impact model reasoning capabilities, or are all CoT approaches equivalent?
To answer this question, we systematically evaluate how the source and validation of rationales affect downstream model performance. We are not only comparing verification strategies, but the different approaches to CoT generation: execution-grounded rationales, LLM-generated rationales (potentially hallucinated), and no rationales at all.
Using the same 25k problem set from our data curation ablation, we generated training data under six distinct conditions and fine-tuned Qwen2.5-Coder-7B models for comparison:
  • Baseline Approaches: (1) Base Model: No fine-tuning; (2) Few-Shot: 3-shot prompting with example problems, no fine-tuning.
  • Alternative CoT Strategies: (3) Q&A Only: Fine-tuned on question-answer pairs with no CoT rationales; (4) LLM-Generated CoT: Fine-tuned on CoT rationales generated purely by an LLM without execution grounding.
  • Our Execution-Grounded Approach: (5) Minimal Execution: Our trace-grounded CoT using minimal verification (1 solution, 1 test); (6) Full Approach: Our complete pipeline with Dual Agreement verification (10 solutions, 25 tests) and trace-grounded rationales.
Table 3 presents the complete results. The ablation reveals a clear hierarchy of components, with three key insights. First, execution grounding is the dominant factor: LLM-generated CoT, despite providing step-by-step rationales, fails because it trains models on plausible but unverified reasoning patterns. This validates our core thesis that the source of rationales—whether grounded in actual execution or generated from statistical patterns—determines whether models internalize verifiable semantics or learn superficial imitation. Second, verification quality compounds this effect: even when using execution-grounded rationales, insufficient candidate diversity degrades performance, as single solution-test pairs cannot filter edge-case failures. This confirms that consensus through Dual Agreement is not merely a scalability mechanism but a necessary correctness guarantee. Third, the step-by-step reasoning itself is essential: removing CoT rationales entirely causes the largest performance drop, demonstrating that models require explicit traces of how outputs arise from inputs, not just input-output correspondences. Prompting alone provides negligible improvement, underscoring that these reasoning capabilities must be internalized through fine-tuning on verified data rather than retrieved at inference time.
Table 3. Impact of verification approach on fine-tuning results. All models fine-tuned on 25k bi-directional data with different verification methods. Tested on Qwen2.5-Coder-7B.
Verification Approach LiveCodeBench CruxEval-O CruxEval-I HumanEval
Base Model (No FT) 46.3 45.3 47.5 62.0
Few-Shot (3 examples) 48.1 47.0 48.8 63.5
Q&A Only (No CoT) 52.7 50.2 51.3 69.0
LLM-as-Judge + LLM CoT 59.3 54.1 55.7 75.5
Our Approach (10S/25T + Trace) 68.2 59.7 61.9 81.5
5.5. Enhancing Instruction-Tuned Models
To confirm that our best-performing bi-directional data provides a specialized signal complementary to general instruction tuning, we fine-tuned existing instruct-tuned models. As shown in Figure 8, this provides a substantial boost, yielding a sharp +39.9 gain on CruxEval Input for granite-3.3-8B-instruct and a +21.5 gain on CruxEval Output for the already strong Qwen2.5-Coder-7B-Instruct, demonstrating the value of our execution-grounded data.
Figure 8. Performance boost from fine-tuning instruct models. Solid bars represent the baseline instruct models; hatched bars show the improvement after fine-tuning with the best-performing 25k bi-directional subset of data.
5.6. Chain-of-Thought Consistency and Quality Analysis
To assess the quality of reasoning generated by our fine-tuned model, we compare CoTs generated by our fine-tuned model Qwen2.5-Coder-7B-FT against the base Qwen2.5-Coder-7B model on CruxEval-O problems. For each problem, we generate CoT from both models and evaluate two key dimensions: consistency between reasoning and final answer, and information content of the reasoning content.
Consistency Analysis. We measure consistency using a composite metric that evaluates reasoning-answer alignment across multiple dimensions: entailment patterns, conceptual overlap, and sequential coherence. Figure 9 (top) shows our model exhibits strong correlation between CoT length and consistency ($R^{2}=0.122$), indicating longer reasoning is more logically coherent. The baseline model shows no such relationship ($R^{2}=0.011$). Moreover, when CoT is consistent with the final answer, our model is more likely to produce correct solutions (AUC = 0.567) compared to the baseline where consistency poorly predicts correctness (AUC = 0.502).
Information Richness. The bottom plot reveals that our model produces substantially more informative reasoning. Using an entropy-based metric that accounts for vocabulary diversity and technical term density, we find the CoTs from our fine-tuned model are 761% more information-rich than those from the baseline. This difference is statistically significant with a very large effect size (Cohen’s $d=7.93$, $p<0.001$), confirming that our training method generates more detailed and semantically meaningful rationales.
These results confirm that verified CoT training fundamentally enhances reasoning quality, producing more consistent and informative chain-of-thought processes that better support final outputs.
Figure 9. Chain-of-Thought quality analysis. Top: CoT-to-outcome consistency vs reasoning length with regression analysis. Bottom: Information content distributions. Verified CoT training shows superior consistency scaling ($R^{2}=0.122$ vs $0.011$) and 761% higher information richness ($d=7.93$, $p<0.001$).
6. Conclusion and Future Work
We introduced a methodology for generating verifiable CoT data for code reasoning, addressing the challenge of logical integrity in synthetic datasets. By grounding every reasoning step in program execution traces, our pipeline eliminates logical hallucinations and produces high-fidelity training data that is correct by construction.
Our systematic evaluation confirms the effectiveness of this approach. We show that verified data quality is more important than quantity and that our novel bi-directional format significantly improves reasoning. These findings hold across different model architectures, where fine-tuning boosted performance by as much as +30.2 points on the CruxEval-O reasoning benchmark, underscoring the fundamental impact of our contribution.
Future work includes extending our language-agnostic pipeline to other languages like C++ and Java and exploring its use in advanced training schemes such as offline reinforcement learning with DPO. To aid the development of more robust language models, we will publicly release our complete synthesis pipeline.
Appendix
This appendix provides concrete examples of the verifiable Chain-of-Thought (CoT) data generated by our synthesis pipeline. The samples are organized to showcase the diversity of our dataset across two key dimensions: reasoning direction and problem difficulty. We include examples for:
  • Reasoning Direction: Forward-only, Backward-only, and Bi-directional conversational formats.
  • Problem Difficulty: Samples rated at levels 2 (Easy), 3 (Medium), and 4 (Hard) based on our quality-scoring mechanism.
Each sample is presented in a structured format that distinguishes the user prompt (the problem context) from the model’s generated response (the verifiable reasoning and final answer).
Evaluation Setup
Our evaluation is conducted using two established, public benchmarks designed to test code reasoning and execution prediction capabilities. We provide direct links to the official benchmark pipelines to ensure full reproducibility of our results. For a fair and transparent comparison across all models, we use the default inference parameters provided by the CruxEval and LiveCodeBench evaluation harnesses without any model-specific tuning.
Experimental Setup
Compute Infrastructure
We perform our experiments on a GPU cluster consisting of Dell XE9680 nodes. Each node has 96 CPU cores with 2 TB of system RAM. Each node is equipped with 8 Nvidia H100 GPUs, each having 80GB of GPU RAM. RHEL 9.4 is installed on the nodes. Each SFT run in our experiments uses 2 nodes (16 GPUs). Evaluation benchmarks are run on a single node using 8 GPUs.
SFT Setup
We use Open Instruct Framework (https://github.com/allenai/open-instruct) for performing SFT. The hyper-parameters used for the training using 16 GPUs are described in table 4.
Batch Size 32
Per Device Batch Size 1
Learning Rate 2e-6
LR Scheduler Linear
Epochs 10
Context Length 8K
Table 4. SFT Hyper-parameters
Data Generation Prompts
This section provides the complete set of prompts used in our data synthesis pipeline. These prompts are referenced in Section 3 (Data Synthesis Pipeline) and are used throughout the hierarchical generation process in Stage A. The prompts are organized by their function in the pipeline, from instruction generation through solution and test case synthesis.
Instruction Generation Prompt
This prompt generates diverse programming instructions for each concept, as described in Section 3.1.
instruction_prompt You are an expert in Python programming and instructional design. Given the concepts and examples below, generate six distinct instructions with complexity level: {difficulty}. Ensure the tasks are as non-overlapping as possible while covering diverse aspects of the concepts. COMPLEXITY GUIDELINES: 1. This {difficulty} difficulty should create {complexity_description}. 2. Solutions should span approximately {expected_lines} of code with rich, intricate logic maximized for hard tasks. INSTRUCTIONS FOR PROBLEM DIVERSITY: 1. Create problems that are fundamentally different in: - Problem domain: Include mathematics (e.g., algebra for equations and transformations, timing & durations for scheduling or sequencing, probabilities for statistical analysis, geometry for spatial computations), finance, data processing, algorithms, text processing, or system design. For hard tasks, prioritize complex mathematical domains to ensure challenging synthesis. - Computational approach: Vary between iterative, recursive, dynamic programming, functional, or object-oriented programming. - Algorithmic complexity: Target specific time complexities (e.g., O(n) for easy, O(n log n) for medium, O(n^2) or higher for hard) appropriate to the difficulty, ensuring hard tasks demand significant computational depth. 2. Before generating, analyze the concepts core principles and identify unique problem-solving strategies that leverage these principles, especially for mathematical domains in hard tasks to maximize complexity and clarity. 3. Instructions may request either a standalone function named solution or a class named Solution with methods; indicate clearly if a class is required (e.g., implement a class’) and specify the primary method name (e.g., compute’) if applicable, otherwise assume compute as the default primary method for classes. Concept: {concept} Description: {description} Examples: {examples} Instructions should contain only the instruction text. Generate your response in the following format: Instruction1: {{}} Instruction2: {{}} Instruction3: {{}} Instruction4: {{}} Instruction5: {{}} Instruction6: {{}}
Signature Generation Prompt
This prompt analyzes instructions and generates appropriate function or class signatures, as described in Section 3.1.
signature_prompt You are a Python programming expert. Given the instruction below, analyze how it should be implemented and provide the best signature skeleton. Follow these STRICT rules to determine the implementation type and format the output: 1. Decide if the instruction requires a standalone function or a class: - Choose a CLASS if the instruction EXPLICITLY says implement a class’, create a class’, or mentions methods like constructor’, build_tree’, etc., using the specified class name (e.g., HuffmanTree’). - Otherwise, default to a standalone FUNCTION named solution’. 2. For a FUNCTION: - Format EXACTLY as: Function: name(param1: type1, param2: type2) -> return_type - Include parameter names, types (infer if not specified), and return type (use unknown if unclear). 3. For a CLASS: - Format EXACTLY as: Class: ClassName; __init__(self, param1: type1) -> return_type; method1(self, param2: type2) -> return_type; ...’ - Use semicolons (;) to separate class name and methods. - Include __init__ with parameters if implied, followed by all required methods. - Specify the primary method (named in instruction or compute if unspecified) for testing. - Use unknown for return types if not inferable. 4. RULES FOR FORMATTING: - Use ONLY spaces (no tabs, newlines, or escaped characters like ’\’). - Use EXACTLY the syntax shown (e.g., __init__’, ’->’, commas between params). - Do NOT add extra punctuation (e.g., colons after parentheses) or quotes around simple types (e.g., use Matrix’, not ’"Matrix"’). - Do NOT deviate from the template---any variation is invalid. - Do NOT include explanations, prose, or multiple class definitions in one block---provide ONLY ONE signature skeleton. - Ensure the response is complete (no truncation) and matches the template EXACTLY. 5. Base your analysis ONLY on the instruction text, inferring types and outputs logically. Instruction: {instruction} Return the signature skeleton INSIDE a code block, following the EXACT format below: ‘‘‘text Function: solution(input1: type1, input2: type2) -> return_type ‘‘‘ or ‘‘‘text Class: ClassName; __init__(self, param1: type1) -> return_type; method1(self, param2: type2) -> return_type ‘‘‘ Examples of CORRECT output: ‘‘‘text Function: solution(freq_list: list[tuple[str, int]]) -> dict[str, str] ‘‘‘ ‘‘‘text Class: HuffmanTree; __init__(self, freq_list: list[tuple[str, int]]) -> unknown; build_tree(self) -> tuple; get_encoding(self) -> dict[str, str] ‘‘‘ Examples of INCORRECT output (DO NOT USE): ‘‘‘text Class: Matrix; **init**(self, data: list[list[int]]): add(self, other: "Matrix") -> "Matrix" ‘‘‘ ‘‘‘text Class: Polynomial; \ __init__(self, coeffs: list[float]) -> None; \ evaluate(self, value: float) -> float ‘‘‘ ‘‘‘text Class: Shape; area(self) -> float; Class: Circle; __init__(self, radius: float) -> None ‘‘‘ Output MUST match the correct examples EXACTLY in format.
Code Generation Prompts
These prompts generate multiple candidate solutions for each instruction. We use separate prompts for function-based and class-based solutions.
Function Code Generation Prompt You are a Python programming expert. Given the instruction and signature details below, generate 5 functionally correct Python code adhering to these constraints: 1. **HIGH PRIORITY**: Implement a standalone function with name ’{function_name}’, inputs ’{input_params}’, and return type ’{return_type}’ EXACTLY as provided. Do NOT deviate from this signature. 2. Write all logic directly within ’{function_name}’---do NOT define nested functions, even for multi-step problems; use variables or steps instead. 3. The function MUST ALWAYS RETURN A VALUE matching ’{return_type}’. 4. Ensure the code is fully modular, self-contained, and does not rely on external code or global variables. 5. Optimize for readability, following Python best practices, with clear variable names and comments where necessary. 6. For hard difficulty, ensure the solution reflects the expected complexity: sophisticated long problems requiring complex algorithms and data structures (8-10 difficulty), spanning approximately 50-100+ lines with a difficulty score of 8-10 on a scale of 1-10. 7. **HIGH PRIORITY**: Generate EXACTLY FIVE distinct implementations, all strictly adhering to the provided signature: - Vary each implementation by: - Computational approach: Use distinct methods like iterative loops, recursion, dynamic programming, list comprehensions, or functional programming (e.g., map/filter/reduce), as appropriate to the instruction and difficulty. - Style: Alternate between verbose, step-by-step logic and concise, optimized solutions; use different commenting styles (e.g., inline vs. block comments). - Variable names: Use unique, meaningful names for variables and parameters in each response. - Complexity: Within the hard level, explore simpler vs. more intricate implementations (e.g., brute force vs. optimized algorithms). - Analyze the instruction to identify multiple viable strategies before generating solutions. - **ENSURE COMPLETENESS**: Each of the five implementations MUST be fully functional, including all required logic as specified in the instruction. Do NOT provide incomplete code (e.g., missing logic); generate all five implementations in full before terminating the response. Instruction: {instruction} Signature Details: - Function Name: {function_name} - Inputs: {input_params} - Return Type: {return_type} Generate EXACTLY FIVE Python code blocks, all adhering to the provided signature, in this format: ‘‘‘python def {function_name}(): # Use the specified function name # Write the solution logic directly here (no nested functions) # Return the final output (MANDATORY) return ... # Replace with actual value ‘‘‘
Class Code Generation Prompt You are a Python programming expert. Given the instruction and signature details below, generate 5 functionally correct Python code implementations adhering to these constraints: 1. HIGH PRIORITY: Implement a class with name ’{class_name}’ and methods as specified in ’{method_signatures}’ (including inputs and return types) EXACTLY as provided. Do NOT deviate from these signature details. 2. Include a constructor ’{constructor_signature}’ ONLY if explicitly provided in the signature details or if the instruction requires initialization of instance variables for the class to function correctly. Otherwise, omit the constructor. 3. Define the class with all necessary methods as specified, avoiding a function template. 4. Each method must be self-contained; each method MUST RETURN A VALUE matching its specified return type. 5. Ensure the code is fully modular, self-contained, and does not rely on external code or global variables. 6. Optimize for readability, following Python best practices, with clear variable names and comments where necessary. 7. For hard difficulty, ensure the solution reflects the expected complexity: sophisticated long problems requiring complex algorithms and data structures (8-10 difficulty), spanning approximately 50-100+ lines with a difficulty score of 8-10 on a scale of 1-10. 8. HIGH PRIORITY: Generate EXACTLY FIVE distinct implementations, all strictly adhering to the provided signature details: - Vary each implementation by: - Computational approach: Use distinct methods like iterative loops, recursion, dynamic programming, list comprehensions, or functional programming (e.g., map/filter/reduce), as appropriate to the instruction and difficulty. - Style: Alternate between verbose, step-by-step logic and concise, optimized solutions; use different commenting styles (e.g., inline vs. block comments). - Variable names: Use unique, meaningful names for variables and parameters in each response. - Complexity: Within the hard level, explore simpler vs. more intricate implementations (e.g., brute force vs. optimized algorithms). - Analyze the instruction to identify multiple viable strategies before generating solutions. - ENSURE COMPLETENESS: Each of the five implementations MUST be fully functional, including all required methods or logic as specified in the instruction. Do NOT provide incomplete code (e.g., missing method bodies or logic); generate all five implementations in full before terminating the response. Instruction: {instruction} Signature Details: - Class Name: {class_name} - Constructor: {constructor_signature} - Methods: {method_signatures} Generate EXACTLY FIVE Python code blocks, all adhering to the provided signature details. Use this format when a constructor is needed: ‘‘‘python class {class_name}: # Use the specified class name def __init__(self, freq_list): # Constructor with specified parameters, only if required # Initialize attributes here pass def build_tree(self): # Specified method # Construct the Huffman tree return ... # Return as required def get_encoding(self): # Specified method # Return the encoding dictionary (MANDATORY) return ... # Replace with actual value ‘‘‘ If no constructor is required, use this simpler format: ‘‘‘python class {class_name}: # Use the specified class name def build_tree(self): # Specified method # Construct the Huffman tree return ... # Return as required def get_encoding(self): # Specified method # Return the encoding dictionary (MANDATORY) return ... # Replace with actual value ‘‘‘
Test Generation Prompts
These prompts generate comprehensive unit test suites for validating the candidate solutions.
Test Scenario Identification Prompt You are an expert in Python testing and requirements analysis. Given the instruction and signature details below, analyze the task and identify a list of up to 10 concise test scenarios to guide test case generation. Each scenario must be a short hint (e.g., Test basic addition’, Test empty input’) to ensure all methods and key behaviors are tested, avoiding excessive detail. Focus on: - Basic functionality of each method or function in the signature. - Key behaviors or operations from the instruction. - Broad coverage of the tasks intent. Return the list in this EXACT format, with no extra text outside the text block: ‘‘‘text Test scenario 1 Test scenario 2 ... ‘‘‘ Task Description: {instruction} Signature Details: {signature_details}
Function Test Generation Prompt You are an expert in Python testing and requirements analysis. Generate up to 10 isolated test cases for the following programming task based on the task description and the provided list of required test scenarios. Follow these CRITICAL GUIDELINES: 1. Each test case must be a standalone Python function (e.g., def test_...():‘), NOT defined within a class, to ensure easy parsing and execution. 2. Each test function must contain EXACTLY ONE assert statement. 3. Every assert statement MUST DIRECTLY call the function with specific inputs and compare its result to an expected value using a direct comparison (e.g., ‘==‘, is‘, in‘, ‘!=‘): - The solution to the task is a standalone function named ’{function_name}’, use assert {function_name}(...) == ...‘ with all inputs packed into the call. - Do NOT: - Use variables or initializations outside the assert (e.g., x = [1, 2]; assert {function_name}(x) == ...‘). - Do NOT Use try-except blocks or check exceptions indirectly (e.g., assert str(e) == ...‘). - Do NOT use vague assertions (e.g., assert == True‘). - Use indirect comparisons (e.g., ‘.equals(...)‘, timing checks). - Rely on external values; pack all necessary logic into the assert statement. 4. Generate up to 10 test cases, each corresponding to one of the required test scenarios provided below, ensuring each test directly calls the function with inputs matching the signature, all within the assert. If fewer than 10 scenarios are provided, generate only that number. 5. Verify that each test aligns with the task requirements, signature details, and the specified test scenario; all inputs must match the provided signature. 6. Ensure every assert statement is complete, specifying a concrete expected output value (e.g., a number, list, or string) and avoiding placeholders (e.g., ’...’). Calculate the exact expected result based on the task description and signature for each test case. Task Description: {instruction} Signature Details: ‘‘‘python {function_signature} ‘‘‘ Required Test Scenarios: {required_tests} Use the template based on the signature (examples show dos and donts): ‘‘‘python # Do this: def test_basic_functionality(): # Test basic scenario assert {function_name}([1, 2, 3], 2) == 42 # Dont do this: def test_basic_functionality_wrong(): # Incorrect: variable outside assert lst = [1, 2, 3] assert {function_name}(lst, 2) == 42 # Dont do this: def test_multi_assert_case(): # Test scenario with multiple independent checks (not preferred) # Test Case 1 assert {function_name}([1, 2], 1) == 10 # Test Case 2 assert {function_name}([3, 4], 1) == 20 ‘‘‘
Class Test Generation Prompt You are an expert in Python testing and requirements analysis. Generate up to 10 isolated test cases for the following programming task based solely on the task description and the provided list of required test scenarios, without seeing the implementation. Follow these CRITICAL GUIDELINES: 1. Each test case must be a standalone Python function (e.g., def test_...():‘), NOT defined within a class, to ensure easy parsing and execution. 2. Each test function must contain EXACTLY ONE assert statement, unless the solution is a class with multiple methods and multiple asserts are needed to call logically connected methods (e.g., setup methods) before the primary method; in such cases, separate each assert with a numbered comment like ‘# Test Case 1‘, ‘# Test Case 2‘, etc., to distinguish them. For connected methods, prefer chaining them within a single assert statement (e.g., ‘{class_name}().setup(...).{primary_method}(...) == ...‘) unless multiple asserts are unavoidable. 3. Every assert statement MUST DIRECTLY call the connected methods with specific inputs and compare its result to an expected value using a direct comparison (e.g., ‘==‘, is‘, in‘, ‘!=‘): - The solution to the task is a class named ’{class_name}’. The primary method to test is ’{primary_method}’. Instantiate it as ‘{class_name}()‘ and call methods directly in the assert; for logically connected methods, chain them within one assert (e.g., assert {class_name}().method1(...).method2(...) == ...‘). Do NOT: - Use variables or class instantiations outside the assert (e.g., c = {class_name}(); assert c.method1(...).method2(...) == ...‘). - Use try-except blocks or check exceptions indirectly (e.g., assert str(e) == ...‘). - Use vague assertions (e.g., assert == True‘). - Use indirect comparisons (e.g., ‘.equals(...)‘, timing checks). - Rely on external values; pack all logic into the assert statement. 4. Generate up to 10 test cases, each corresponding to one of the required test scenarios provided below, ensuring each test directly calls the relevant method(s) with inputs matching their signature, all within the assert. If fewer than 10 scenarios are provided, generate only that number. 5. Verify that each test aligns with the task requirements, signature details, and the specified test scenario; all inputs must match the method signatures. 6. Ensure every assert statement is complete, specifying a concrete expected output value (e.g., a number, list, or string) and avoiding placeholders (e.g., ’...’). Calculate the exact expected result based on the task description and signature for each test case. Task Description: {instruction} Signature Details: ‘‘‘python Class: {class_name} Class Methods: {method_signatures} Primary Method: {primary_method} ‘‘‘ Required Test Scenarios: {required_tests} Generate test cases in this format, with each test in its own standalone function, using ONLY direct calls in asserts with complete expected values, packing all logic into the assert. Use the template based on the signature (examples show dos and donts): - For class-based solutions: ‘‘‘python # Do this: def test_basic_functionality(): # Test basic scenario assert {class_name}().{primary_method}([1, 2, 3]) == 42 # Do this for logically connected methods, ensuring instantiation and calls are in one assert: def test_connected_methods(): # Test scenario where object instantiation and connected method calls are all in one assert statement assert {class_name}().setup([1, 2]).{primary_method}(3) == 42 # Dont do this: def test_basic_functionality_wrong(): # Incorrect: multiple asserts for class without logical connection assert {class_name}().{primary_method}([1, 2]) == 10 assert {class_name}().{primary_method}([3, 4]) == 20 # Dont do this: def test_setup_wrong(): # Incorrect: setup outside assert obj = {class_name}() obj.setup([1, 2]) assert obj.{primary_method}(3) == 42 ‘‘‘
All custom analysis and evaluation scripts used to generate the results and figures presented in this paper, beyond the benchmark pipelines themselves, are included as part of the supplementary material.
Appendix: Verifiable CoT Data Samples
This appendix provides concrete examples of the verifiable Chain-of-Thought (CoT) data generated by our synthesis pipeline. The samples are organized to showcase the diversity of our dataset across two key dimensions: reasoning direction and problem difficulty. We include examples for:
  • Reasoning Direction: Forward-only, Backward-only, and Bi-directional conversational formats.
  • Problem Difficulty: Samples rated at levels 2 (Easy), 3 (Medium), and 4 (Hard) based on our quality-scoring mechanism.
Each sample is presented in a structured format that distinguishes the user prompt (the problem context) from the model’s generated response (the verifiable reasoning and final answer).
Appendix Sample: Forward Reasoning (Difficulty 2)
User Prompt (Forward) The ‘solution‘ function takes a list of events with associated probabilities, normalizes them if they do not sum to 1, and returns the cumulative distribution function (CDF) and the probability density function (PDF). Here’s the code: from functools import reduce def solution(events: list[tuple[str, float]]) -> tuple[list[float], list[float]]: # Sort events by probability sorted_events = sorted(events, key=lambda x: x[1]) # Extract probabilities probabilities = list(map(lambda x: x[1], sorted_events)) # Normalize if necessary total_prob = sum(probabilities) if total_prob != 1: probabilities = [p / total_prob for p in probabilities] # Calculate PDF pdf = probabilities # Calculate CDF cdf = [] cumulative_prob = 0 for prob in pdf: cumulative_prob += prob cdf.append(cumulative_prob) return cdf, pdf What output does the function return when given the input [(’a’, 0.5), (’b’, 0.5)]?
Assistant Response ### Understand The goal is to determine the output of the function solution when given the input events = [(’a’, 0.5), (’b’, 0.5)]‘. The functions purpose is to compute two lists: the probability density function (PDF) and the cumulative distribution function (CDF) of the input probabilities. The function sorts the events by their probabilities, extracts the probabilities, normalizes them if necessary, and then calculates the PDF and CDF. Key components: - Sorting the events by probability. - Extracting probabilities and normalizing them if they dont sum to 1. - Calculating the PDF (normalized probabilities). - Calculating the CDF by accumulating the probabilities. ### Plan To determine the output, I will: 1. Analyze the input and the initial operations (sorting and extracting probabilities). 2. Check if normalization is needed and perform it if necessary. 3. Calculate the PDF using the normalized probabilities. 4. Calculate the CDF by accumulating the PDF values. 5. Verify the final output by considering the logical flow and ensuring all steps are consistent. ### Execute 1. **Input and Sorting**: The input is events = [(’a’, 0.5), (’b’, 0.5)]‘. The sorting step sorts the events by the second element (probability), but since both probabilities are equal, the order remains ‘(’a’, 0.5), (’b’, 0.5)‘. 2. **Extracting Probabilities**: The probabilities extracted are ‘[0.5, 0.5]‘. 3. **Normalization Check**: The sum of probabilities is ‘0.5 + 0.5 = 1.0‘. Since the total is already 1, no normalization is needed, and the probabilities remain ‘[0.5, 0.5]‘. 4. **PDF Calculation**: The PDF is directly the list of probabilities, ‘[0.5, 0.5]‘. 5. **CDF Calculation**: - Initialize cumulative_prob = 0‘. - For the first probability ‘0.5‘, cumulative_prob = 0 + 0.5 = 0.5‘. Append ‘0.5‘ to cdf‘. - For the second probability ‘0.5‘, cumulative_prob = 0.5 + 0.5 = 1.0‘. Append ‘1.0‘ to cdf‘. - The resulting CDF is ‘[0.5, 1.0]‘. 6. **Final Output**: The function returns the tuple ‘(cdf, pdf)‘, which is ‘([0.5, 1.0], [0.5, 0.5])‘. ### Reflect The reasoning process involved analyzing each step of the function, ensuring that the probabilities were correctly extracted, checked for normalization, and used to compute the PDF and CDF. The calculations were straightforward, and the logic of accumulating probabilities for the CDF was correctly applied. The final output aligns with the expected behavior of the function, confirming that the reasoning was thorough and accurate. Predicted Output: ([0.5, 1.0], [0.5, 0.5])
Appendix Sample: Forward Reasoning (Difficulty 3)
User Prompt (Forward) The ‘solution‘ function approximates a solution to the Traveling Salesman Problem by using a greedy algorithm to find a route that visits all given locations exactly once, prioritizing closer locations (within 10 units) to the current location, and returns to the starting point. Here’s the code: def solution(locations): from math import sqrt def distance(p1, p2): return sqrt((p1[0] - p2[0]) ** 2 + (p1[1] - p2[1]) ** 2) if len(locations) < 2: return locations # Start from the first location start = locations[0] remaining = locations[1:] route = [start] while remaining: # Calculate distances from the current location current = route[-1] # Prioritize locations within 10 units using ternary operation prioritized = sorted( remaining, key=lambda loc: (distance(current, loc) <= 10, distance(current, loc)) ) # Select the closest location next_location = prioritized[0] route.append(next_location) remaining.remove(next_location) # Return to the starting point route.append(start) return route What does the function return when given the input [(0, 0), (1, 1), (2, 2), (3, 3)]?
Assistant Response ### Understand The goal is to determine the output of the solution function when given the input ‘[(0, 0), (1, 1), (2, 2), (3, 3)]‘. The function is designed to construct a route from a list of locations, starting from the first location, and then iteratively selecting the next location based on proximity, prioritizing those within 10 units. The function returns to the starting point after visiting all other locations. Key components include the distance function, the sorting mechanism for prioritizing locations, and the loop that builds the route. ### Plan 1. **Initialization**: Understand the initial setup of variables like start‘, remaining‘, and route‘. 2. **Loop Execution**: Analyze the loops behavior, focusing on how it selects the next location based on distance and prioritization. 3. **Return to Start**: Consider the final step where the route returns to the starting point. 4. **Verify**: Ensure the reasoning aligns with the expected behavior of the function. ### Execute 1. **Initialization**: - The function begins by checking if there are fewer than two locations, in which case it returns the input as is. This is not applicable here since there are four locations. - start is set to the first location ‘(0, 0)‘. - remaining is initialized with the rest of the locations ‘[(1, 1), (2, 2), (3, 3)]‘. - route starts with ‘[start]‘, i.e., ‘[(0, 0)]‘. 2. **Loop Execution**: - The loop continues as long as there are locations in remaining‘. - For each iteration, the current location is the last in route‘. - The remaining locations are sorted based on whether they are within 10 units and then by distance. - The closest location is selected and added to route‘, then removed from remaining‘. **Iteration Details**: - **First Iteration**: - current is ‘(0, 0)‘. - Distances to remaining locations: - ‘(1, 1)‘: $\(\sqrt${2}\) - ‘(2, 2)‘: $\(\sqrt${8}\) - ‘(3, 3)‘: $\(\sqrt${18}\) - All are within 10 units, so they are sorted by distance: ‘[(1, 1), (2, 2), (3, 3)]‘. - next_location is ‘(1, 1)‘, added to route‘, and removed from remaining‘. - **Second Iteration**: - current is ‘(1, 1)‘. - Distances to remaining locations: - ‘(2, 2)‘: $\(\sqrt${2}\) - ‘(3, 3)‘: $\(\sqrt${8}\) - Both are within 10 units, sorted by distance: ‘[(2, 2), (3, 3)]‘. - next_location is ‘(2, 2)‘, added to route‘, and removed from remaining‘. - **Third Iteration**: - current is ‘(2, 2)‘. - Distance to remaining location ‘(3, 3)‘ is $\(\sqrt${2}\). - next_location is ‘(3, 3)‘, added to route‘, and removed from remaining‘. 3. **Return to Start**: - After the loop, the route returns to the starting point ‘(0, 0)‘ by appending it to route‘. 4. **Final Route**: - The constructed route is ‘[(0, 0), (1, 1), (2, 2), (3, 3), (0, 0)]‘. ### Reflect The reasoning process aligns with the functions logic, ensuring that each step is consistent with the codes behavior. The function correctly constructs a route based on proximity and returns to the start, confirming the expected output. The approach considered the sorting mechanism and loop behavior, ensuring a thorough understanding of the functions execution. The final output is consistent with the expected behavior and the trace provided. Predicted Output: [(0, 0), (1, 1), (2, 2), (3, 3), (0, 0)]
Appendix Sample: Forward Reasoning (Difficulty 4)
User Prompt (Forward) The ‘solution‘ function identifies the pair of tuples with the smallest Euclidean distance from a sorted list of coordinate tuples using a divide-and-conquer approach, ensuring efficient performance with a time complexity of O(n log n). Here’s the code: def solution(coordinates): # Helper function to calculate Euclidean distance def euclidean_distance(point1, point2): return ((point1[0] - point2[0]) ** 2 + (point1[1] - point2[1]) ** 2) ** 0.5 # Recursive function to find the closest pair def closest_pair(coords): if len(coords) <= 3: return brute_force(coords) mid = len(coords) // 2 left_half = coords[:mid] right_half = coords[mid:] # Find closest pairs in each half left_closest = closest_pair(left_half) right_closest = closest_pair(right_half) # Determine the smaller distance min_distance = min(euclidean_distance(*left_closest), euclidean_distance(*right_closest)) closest_pair = left_closest if euclidean_distance(*left_closest) < euclidean_distance(*right_closest) else right_closest # Check the strip around the midpoint strip = [point for point in coords if abs(point[0] - coords[mid][0]) < min_distance] strip.sort(key=lambda x: x[1]) # Sort strip by y-coordinate for i in range(len(strip)): for j in range(i + 1, len(strip)): if (strip[j][1] - strip[i][1]) >= min_distance: break distance = euclidean_distance(strip[i], strip[j]) if distance < min_distance: min_distance = distance closest_pair = (strip[i], strip[j]) return closest_pair # Brute force approach for small number of points def brute_force(points): min_distance = float(’inf’) closest = None for i in range(len(points)): for j in range(i + 1, len(points)): distance = euclidean_distance(points[i], points[j]) if distance < min_distance: min_distance = distance closest = (points[i], points[j]) return closest # Sort coordinates by x-coordinate sorted_coords = sorted(coordinates, key=lambda x: x[0]) return closest_pair(sorted_coords) What output does the ‘solution‘ function return when given the input ‘[(1, 2), (3, 4)]‘?
Assistant Response ### Understand The solution function is designed to find the closest pair of points from a list of 2D coordinates using Euclidean distance. The function employs a recursive approach to solve this problem efficiently, leveraging a divide-and-conquer strategy. The key components of the function include: 1. **Euclidean Distance Calculation**: A helper function euclidean_distance computes the distance between two points. 2. **Recursive Closest Pair Function**: The closest_pair function recursively divides the list of points and finds the closest pair in each half. It also considers points near the dividing line to ensure the closest pair is found. 3. **Brute Force Method**: For small subsets of points (3 or fewer), the brute_force function is used to find the closest pair by checking all possible pairs. 4. **Sorting**: The input coordinates are sorted by their x-coordinates before processing. The functions goal is to return the pair of points with the smallest Euclidean distance. ### Plan To determine the output for the input ‘[(1, 2), (3, 4)]‘, we will: 1. **Analyze Initialization**: Understand how the input is processed initially, particularly the sorting step. 2. **Trace Recursive Logic**: Follow the recursive logic of closest_pair to see how it handles the input. 3. **Evaluate Base Case**: Since the input has only two points, the base case of the recursion will be reached quickly. 4. **Compute Distances**: Calculate the Euclidean distance between the two points to verify the result. 5. **Consider Edge Cases**: Although not directly applicable here, consider if the logic holds for minimal inputs. ### Execute 1. **Initialization**: The input ‘[(1, 2), (3, 4)]‘ is sorted by x-coordinates, resulting in sorted_coords = [(1, 2), (3, 4)]‘. 2. **Recursive Logic**: - The closest_pair function is called with sorted_coords‘. - Since the length of coords is 2, the function does not split further and directly proceeds to the brute_force function. 3. **Base Case Evaluation**: - In brute_force‘, the function iterates over all pairs of points. - For ‘[(1, 2), (3, 4)]‘, there is only one pair to consider: ‘(1, 2)‘ and ‘(3, 4)‘. 4. **Distance Calculation**: - Compute the Euclidean distance between ‘(1, 2)‘ and ‘(3, 4)‘: \[ $\text${distance} = $\sqrt${(3 - 1)^2 + (4 - 2)^2} = $\sqrt${4 + 4} = $\sqrt${8} = $\sqrt{2}$ \] - Since this is the only pair, it is the closest pair by default. 5. **Return Result**: - The function returns the pair ‘((1, 2), (3, 4))‘ as the closest pair. ### Reflect The reasoning aligns with the functions logic, as the input size is small enough to directly use the brute force method. The calculation of the Euclidean distance confirms that the pair ‘(1, 2)‘ and ‘(3, 4)‘ is indeed the closest pair. The functions design ensures that even with minimal inputs, the correct pair is returned. The approach is robust and generalizes well to larger inputs by using the divide-and-conquer strategy. Predicted Output: ((1, 2), (3, 4))
Appendix Sample: Forward Reasoning (Difficulty 4)
User Prompt (Forward) The ‘solution‘ function calculates the probability of an event occurring across a sequence of independent trials using binary partitioning and the inclusion-exclusion principle, efficiently handling up to 1,000 trials with memoization to ensure high precision. Here’s the code: def solution(probabilities): # Helper function to calculate the combined probability using inclusion-exclusion def combine_probs(p1, p2): return p1 + p2 - (p1 * p2) # Recursive approach with memoization to avoid redundant calculations memo = {} def recursive_prob(start, end): if (start, end) in memo: return memo[(start, end)] if end - start == 1: return probabilities[start] mid = (start + end) // 2 prob_left = recursive_prob(start, mid) prob_right = recursive_prob(mid, end) combined = combine_probs(prob_left, prob_right) memo[(start, end)] = combined return combined return recursive_prob(0, len(probabilities)) When I pass a list of probabilities to the ‘solution‘ function, does it return a combined probability of 1.0, considering it uses the inclusion-exclusion principle to calculate the probability?
Assistant Response ### Understand The solution function is designed to compute a combined probability from a list of individual probabilities using the inclusion-exclusion principle. The function includes a helper function combine_probs that combines two probabilities using the formula \( p1 + p2 - (p1 $\times$ p2) \). The main function uses a recursive approach with memoization to efficiently calculate the combined probability for the entire list. The recursion splits the list into halves until it reaches base cases of single probabilities, then combines them back up using the combine_probs function. The memoization helps avoid redundant calculations by storing results of subproblems. ### Plan To determine if the function returns a combined probability of 1.0, we need to: 1. Understand how the inclusion-exclusion principle is applied through the combine_probs function. 2. Analyze the recursive structure and how it processes the list of probabilities. 3. Consider edge cases, such as when the list contains probabilities that sum to 1 or when probabilities are 0 or 1. 4. Verify if the recursive combination of probabilities can logically result in 1.0. ### Execute 1. **Inclusion-Exclusion Principle**: The combine_probs function combines two probabilities \( p1 \) and \( p2 \) using the formula \( p1 + p2 - (p1 $\times$ p2) \). This formula accounts for the overlap between two events, ensuring that the combined probability does not exceed 1. 2. **Recursive Structure**: The function recursive_prob divides the list into smaller segments, computes the combined probability for each segment, and then combines these results. The base case is when the segment contains a single probability, which is returned directly. 3. **Edge Cases**: - If all probabilities are 0, the combined probability will be 0. - If all probabilities are 1, the combined probability will be 1. - If the list contains a mix of 0s and 1s, the result will depend on their arrangement and combination. 4. **Logical Combination to 1.0**: For the function to return 1.0, the combination of probabilities must account for all possible overlaps such that the final result is 1.0. This can happen if the probabilities are structured such that their combined effect covers the entire probability space. 5. **Mathematical Consideration**: The inclusion-exclusion principle ensures that as more probabilities are combined, the overlap is subtracted, preventing the result from exceeding 1.0. Given enough probabilities and their overlaps, it is possible for the combined probability to reach 1.0, especially if the probabilities are structured to cover all possible outcomes. ### Reflect The functions logic, using the inclusion-exclusion principle and recursive combination, can indeed result in a combined probability of 1.0, particularly if the input probabilities are structured to cover the entire probability space. The recursive approach with memoization ensures efficient computation, and the inclusion-exclusion principle prevents the result from exceeding 1.0. The trace confirms that the function returns 1.0 for the given input, aligning with the logical deduction. Predicted Output: True
Appendix Sample: Forward Reasoning (Difficulty 4)
User Prompt (Forward) The function schedules tasks by sorting them by deadline and using dynamic programming to determine the minimum start time for each task, then reconstructs and returns the schedule with task details including start times. Here’s the code: def solution(tasks): # Sort tasks by their deadline tasks.sort(key=lambda x: x[2]) n = len(tasks) # Initialize a DP table to store the minimum start time for each task dp = [[float(’inf’)] * (n + 1) for _ in range(n + 1)] dp[0][0] = 0 # Fill the DP table for i in range(1, n + 1): task_name, duration, deadline = tasks[i - 1] for j in range(i + 1): # If we don’t take the current task dp[i][j] = dp[i - 1][j] # If we take the current task if j > 0: start_time = max(dp[i - 1][j - 1], deadline - duration) dp[i][j] = min(dp[i][j], start_time + duration) # Reconstruct the solution result = [] time = dp[n][n] for i in range(n, 0, -1): if dp[i][n] != dp[i - 1][n]: task_name, duration, deadline = tasks[i - 1] start_time = max(dp[i - 1][n - 1], deadline - duration) result.append((task_name, duration, deadline, start_time)) time = start_time n -= 1 return result[::-1] When I provide the input ‘[("task1", 0, 0), ("task2", 2147483647, 2147483647)]‘ to the ‘solution‘ function, will it return ‘[(’task1’, 0, 0, 0), (’task2’, 2147483647, 2147483647, 0)]‘ as expected?
Assistant Response ### Understand The goal is to determine if the solution function, when given the input ‘[("task1", 0, 0), ("task2", 2147483647, 2147483647)]‘, will return ‘[(’task1’, 0, 0, 0), (’task2’, 2147483647, 2147483647, 0)]‘. The function appears to be solving a scheduling problem using dynamic programming (DP). It sorts tasks by their deadlines and calculates the earliest possible start time for each task, ensuring no task starts after its deadline. The DP table dp is used to store the minimum start time for scheduling a subset of tasks. The function then reconstructs the schedule to provide the start time for each task. ### Plan 1. **Initialization and Sorting**: Verify that the tasks are sorted by their deadlines. 2. **DP Table Setup**: Understand the initialization of the DP table and its purpose. 3. **DP Table Filling**: Analyze how the DP table is filled, focusing on the logic for deciding whether to include a task and calculating the start time. 4. **Reconstruction**: Examine how the solution is reconstructed from the DP table. 5. **Validation**: Check if the reconstructed solution matches the expected output. ### Execute 1. **Initialization and Sorting**: - The input tasks are sorted by their deadlines: ‘[("task1", 0, 0), ("task2", 2147483647, 2147483647)]‘. Since "task1" has an earlier deadline, the order remains unchanged. 2. **DP Table Setup**: - A DP table dp of size ‘(n+1) x (n+1)‘ is initialized with inf‘, except dp[0][0] = 0‘. This setup is for storing the minimum start time for scheduling j tasks out of the first i tasks. 3. **DP Table Filling**: - For each task, the table is filled considering two scenarios: including the task or not. - **Task 1 ("task1", 0, 0)**: - For j = 0‘: dp[1][0] = dp[0][0] = 0‘. - For j = 1‘: Calculate start_time = max(dp[0][0], 0 - 0) = 0‘. Update dp[1][1] = min(inf, 0 + 0) = 0‘. - **Task 2 ("task2", 2147483647, 2147483647)**: - For j = 0‘: dp[2][0] = dp[1][0] = 0‘. - For j = 1‘: Calculate start_time = max(dp[1][0], 2147483647 - 2147483647) = 0‘. Update dp[2][1] = min(0, 0 + 2147483647) = 0‘. - For j = 2‘: Calculate start_time = max(dp[1][1], 2147483647 - 2147483647) = 0‘. Update dp[2][2] = min(inf, 0 + 2147483647) = 2147483647‘. 4. **Reconstruction**: - Start from the last task and check if including it changes the DP value. - For i = 2‘ ("task2"): dp[2][2] != dp[1][2]‘, so include it. Calculate start_time = max(dp[1][1], 2147483647 - 2147483647) = 0‘. Add ‘(’task2’, 2147483647, 2147483647, 0)‘. - Update time = 0‘, n = 1‘. - For i = 1‘ ("task1"): dp[1][1] != dp[0][1]‘, so include it. Calculate start_time = max(dp[0][0], 0 - 0) = 0‘. Add ‘(’task1’, 0, 0, 0)‘. 5. **Validation**: - The reconstructed result is ‘[(’task2’, 2147483647, 2147483647, 0), (’task1’, 0, 0, 0)]‘. - Reverse it to get ‘[(’task1’, 0, 0, 0), (’task2’, 2147483647, 2147483647, 0)]‘, which matches the expected output. ### Reflect The reasoning process involved understanding the DP approach used for scheduling tasks and verifying each step of the logic. The reconstruction step was crucial in ensuring the output matched the expected result. The chosen method was robust, as it considered both the DP table filling and the reconstruction process. The final output aligns with the expected result, confirming the correctness of the functions logic. Predicted Output: [(’task1’, 0, 0, 0), (’task2’, 2147483647, 2147483647, 0)]
Appendix Sample: Backward Reasoning (Difficulty 4)
User Prompt (Backward) The function ‘solution‘ calculates and returns the mean, median, and mode of a non-empty list of integers, ensuring input validity through assertions. Here’s the code: def solution(numbers: list[int]) -> dict[str, float]: # Assertions to ensure valid input assert numbers, "Input list cannot be empty." assert all(isinstance(item, int) for item in numbers), "List must contain only integers." # Mean calculation total = sum(numbers) length = len(numbers) mean_result = total / length # Median calculation sorted_nums = sorted(numbers) half = length // 2 if length % 2 == 0: median_result = (sorted_nums[half - 1] + sorted_nums[half]) / 2 else: median_result = sorted_nums[half] # Mode calculation using a dictionary to count occurrences count_dict = {} for num in numbers: count_dict[num] = count_dict.get(num, 0) + 1 max_count = max(count_dict.values()) mode_result = min(key for key, value in count_dict.items() if value == max_count) # Return the computed statistics return {’mean’: mean_result, ’median’: median_result, ’mode’: float(mode_result)} What input would produce the output ’mean’: 2.0, ’median’: 2, ’mode’: 1 when passed to the function?
Assistant Response ### Understand The goal is to deduce the input(s) that produce the output ‘{’mean’: 2.0, median’: 2, mode’: 1}‘ when passed to the solution function. The function calculates three statistical measures: mean, median, and mode from a list of integers. The mean is the average of the numbers, the median is the middle value when sorted, and the mode is the most frequent number (with ties broken by the smallest number). The function includes assertions to ensure the input is a non-empty list of integers. ### Plan To deduce the input(s), I will reverse engineer the calculations for mean, median, and mode: 1. **Mean**: The mean is calculated as the sum of the numbers divided by the count. For a mean of 2.0, the sum of the numbers divided by their count must equal 2.0. 2. **Median**: The median is the middle value in a sorted list. For a median of 2, the middle value (or average of two middle values) must be 2. 3. **Mode**: The mode is the most frequent number. For a mode of 1, the number 1 must appear more frequently than any other number. I will explore different list configurations that satisfy these conditions, considering both odd and even lengths of the list. I will also ensure that the mode condition is met by having 1 as the most frequent number. ### Execute 1. **Mean Calculation**: - Let the list be ‘[a1, a2, ..., an]‘. - The mean condition is: \((a1 + a2 + ... + an) / n = 2.0\). - Therefore, \(a1 + a2 + ... + an = 2n\). 2. **Median Calculation**: - For an odd-length list, the median is the middle element. - For an even-length list, the median is the average of the two middle elements. - The median must be 2. 3. **Mode Calculation**: - The number 1 must appear more frequently than any other number. **Case 1: Odd-length list (n = 3)** - Mean: \(a1 + a2 + a3 = 6\). - Median: The middle element must be 2. - Mode: 1 must appear more frequently than any other number. Possible list: ‘[1, 2, 3]‘ - Sum: \(1 + 2 + 3 = 6\), Mean: \(6/3 = 2.0\). - Sorted: ‘[1, 2, 3]‘, Median: 2. - Mode: All numbers appear once, but 1 is the smallest, so mode is 1. **Case 2: Even-length list (n = 4)** - Mean: \(a1 + a2 + a3 + a4 = 8\). - Median: The average of the two middle elements must be 2. - Mode: 1 must appear more frequently than any other number. Possible list: ‘[1, 1, 2, 4]‘ - Sum: \(1 + 1 + 2 + 4 = 8\), Mean: \(8/4 = 2.0\). - Sorted: ‘[1, 1, 2, 4]‘, Median: \((1 + 2)/2 = 1.5\) (not valid). **Case 3: Odd-length list (n = 5)** - Mean: \(a1 + a2 + a3 + a4 + a5 = 10\). - Median: The middle element must be 2. - Mode: 1 must appear more frequently than any other number. Possible list: ‘[1, 1, 2, 3, 3]‘ - Sum: \(1 + 1 + 2 + 3 + 3 = 10\), Mean: \(10/5 = 2.0\). - Sorted: ‘[1, 1, 2, 3, 3]‘, Median: 2. - Mode: 1 appears twice, 3 appears twice, but 1 is the smallest, so mode is 1. ### Reflect The deduced inputs that satisfy the conditions are ‘[1, 2, 3]‘ and ‘[1, 1, 2, 3, 3]‘. Both produce the desired output. The given input ‘[1, 2, 3]‘ matches one of the deduced inputs, confirming the correctness of the deduction. The approach was exhaustive, considering both odd and even lengths, and ensuring all conditions for mean, median, and mode were met. Predicted Input: Plausible input 1: [1, 2, 3] Plausible input 2: [1, 1, 2, 3, 3]
Appendix Sample: Backward Reasoning (Difficulty 4)
User Prompt (Backward) The ‘invert_matrix‘ method in the ‘MatrixOperations‘ class calculates and returns the inverse of a given square matrix (up to 3x3), ensuring the matrix is square and non-singular with assert statements. Here’s the code: class MatrixOperations: def invert_matrix(self, matrix: list[list[float]]) -> list[list[float]]: n = len(matrix) assert all(len(row) == n for row in matrix), "Matrix must be square" det = self._recursive_determinant(matrix) assert det != 0, "Matrix is singular and cannot be inverted" cofactors = [[self._cofactor(matrix, i, j) for j in range(n)] for i in range(n)] adjugate = [[cofactors[j][i] for j in range(n)] for i in range(n)] inverse = [[adjugate[i][j] / det for j in range(n)] for i in range(n)] return inverse def _recursive_determinant(self, matrix: list[list[float]]) -> float: n = len(matrix) if n == 1: return matrix[0][0] if n == 2: return matrix[0][0] * matrix[1][1] - matrix[0][1] * matrix[1][0] det = 0 for c in range(n): det += ((-1) ** c) * matrix[0][c] * self._recursive_determinant(self._minor(matrix, 0, c)) return det def _minor(self, matrix: list[list[float]], i: int, j: int) -> list[list[float]]: return [row[:j] + row[j+1:] for row in (matrix[:i] + matrix[i+1:])] def _cofactor(self, matrix: list[list[float]], i: int, j: int) -> float: minor = self._minor(matrix, i, j) return ((-1) ** (i + j)) * self._recursive_determinant(minor) What input matrix would produce the inverse ‘[[0.6, -0.7], [-0.2, 0.4]]‘ when passed to the ‘invert_matrix‘ function?
[…]
Appendix Sample: Bi-Directional Conversation (Difficulty 4)
User Prompt (Forward) The ‘solution‘ function sorts a list of DNA sequences by their GC-content in descending order, using lexicographical order as a tiebreaker, and includes error handling for invalid DNA sequences containing characters other than ’A’, ’T’, ’C’, ’G’. Here’s the code: def solution(dna_sequences): # Calculate GC-content using list comprehension def gc_content(seq): return (seq.count(’G’) + seq.count(’C’)) / len(seq) # Validate sequences and prepare for sorting valid_sequences = [ (gc_content(seq), seq) for seq in dna_sequences if all(nucleotide in ’ATCG’ for nucleotide in seq) ] # Raise error if any invalid sequence is found if len(valid_sequences) != len(dna_sequences): raise ValueError("Invalid DNA sequence detected.") # Sort sequences by GC-content and lexicographically valid_sequences.sort(key=lambda x: (-x[0], x[1])) # Return sorted sequences return [seq for _, seq in valid_sequences] When we pass a list of DNA sequences where each sequence is composed of a varying number of ’G’s followed by ’A’s, does the function return the sequences sorted by their GC-content from lowest to highest?
Assistant Response ### Understand The task is to determine if the given Python function solution sorts a list of DNA sequences by their GC-content from lowest to highest. Each sequence is composed of a varying number of Gs followed by As. The function calculates the GC-content for each sequence, validates the sequences, and sorts them based on GC-content and lexicographically. Key components include the gc_content function, the validation of sequences, and the sorting mechanism. ### Plan 1. **Analyze the gc_content function**: Understand how it calculates the GC-content for a sequence. 2. **Examine sequence validation**: Ensure all sequences are valid DNA sequences. 3. **Understand the sorting logic**: Determine how sequences are sorted by GC-content and lexicographically. 4. **Evaluate the overall effect**: Assess if the function sorts sequences from lowest to highest GC-content. 5. **Consider edge cases**: Check if sequences with the same GC-content are handled correctly. ### Execute 1. **GC-content Calculation**: - The gc_content function calculates the GC-content as the ratio of Gs and Cs to the total length of the sequence. Since sequences only contain Gs and As, Cs are absent, and the GC-content is simply the proportion of Gs. 2. **Sequence Validation**: - The function checks if each sequence contains only A’, T’, C’, and G’. Given the problem constraints, all sequences are valid. 3. **Sorting Logic**: - The sequences are sorted using valid_sequences.sort(key=lambda x: (-x[0], x[1]))‘. This sorts primarily by GC-content in descending order (‘-x[0]‘) and secondarily by lexicographical order (‘x[1]‘). 4. **Effect of Sorting**: - The sorting order is from highest to lowest GC-content due to the negative sign in ‘-x[0]‘. This means sequences with more Gs (higher GC-content) appear first. 5. **Edge Cases**: - If sequences have the same GC-content, they are sorted lexicographically. Since G < A’, sequences with more Gs will appear first among those with the same GC-content. ### Reflect The function does not sort sequences from lowest to highest GC-content; it sorts them from highest to lowest. The use of ‘-x[0]‘ in the sorting key indicates descending order. This conclusion aligns with the functions logic and the observed behavior in the execution trace. The reasoning process considered the functions structure, calculations, and sorting mechanism, ensuring a thorough understanding. Predicted Output: No, the function returns the sequences sorted by their GC-content from highest to lowest.
User Prompt (Backward) What kind of input list would result in the function returning a list of DNA sequences sorted from those with 0
Assistant Response ### Understand The goal is to determine what kind of input list would cause the function solution to return a list of DNA sequences sorted from those with 0% GC-content to those with 100% GC-content. The function sorts them primarily by GC-content in descending order. The final output is a list of sequences sorted in descending order of GC-content, which is the opposite of the desired order (ascending). Therefore, we need to reverse this logic to deduce the input that would result in the desired output. ### Plan 1. **Reverse the Sorting Logic**: Since the function sorts sequences by descending GC-content, we need to consider how the input could be structured so that when sorted in descending order, the sequences appear in ascending order of GC-content. 2. **Identify GC-content Calculation**: Understand how GC-content is calculated and how it influences sorting. 3. **Consider Edge Cases**: Explore sequences with extreme GC-content values (0% and 100%) and how they affect sorting. ### Execute 1. **Reverse the Sorting Logic**: - The function sorts sequences by ‘(-x[0], x[1])‘, meaning it first sorts by GC-content in descending order. To achieve ascending order of GC-content, the input should be structured such that the descending sort results in the desired order. 2. **Identify GC-content Calculation**: - GC-content is ‘(seq.count(’G’) + seq.count(’C’)) / len(seq)‘. Sequences with more G and C have higher GC-content. 3. **Deduce Input**: - To achieve the desired output, input sequences should be structured such that when sorted by descending GC-content, they naturally appear in ascending order. - Input: ‘[’AAAA’, AACG’, ACGT’, CGCG’, GGGG’]‘ - This input results in GC-content values of 0%, 50%, 50%, 100%, and 100%, respectively. When sorted by descending GC-content, the order is ‘[’GGGG’, CGCG’, ACGT’, AACG’, AAAA’]‘, which is the reverse of the desired order. ### Reflect The deduced input is structured to ensure that when sorted by descending GC-content, the sequences appear in ascending order of GC-content. This approach considers the functions sorting logic and the calculation of GC-content. Predicted Input: Plausible input 1: [’AAAA’, ’AACG’, ’ACGT’, ’CGCG’, ’GGGG’]
Appendix Sample: Bi-Directional Conversation (Difficulty 4)
User Prompt (Forward) The ‘solve_polynomial‘ function finds all real roots of a polynomial equation using Newton’s method, handling polynomials up to degree 10, with mechanisms to estimate the derivative and manage non-convergence cases. Here’s the code: def solve_polynomial(coefficients: list[float]) -> list[float]: """ Solves polynomial equations using a recursive implementation of Newton’s method. """ from math import isclose def evaluate_polynomial(coeffs, x): return sum(c * (x ** i) for i, c in enumerate(coeffs)) def evaluate_derivative(coeffs, x): return sum(i * c * (x ** (i - 1)) for i, c in enumerate(coeffs) if i != 0) def newtons_method_recursive(coeffs, x, tolerance=1e-7, max_iterations=1000, iteration=0): if iteration >= max_iterations: return None f_x = evaluate_polynomial(coeffs, x) f_prime_x = evaluate_derivative(coeffs, x) if isclose(f_prime_x, 0, abs_tol=tolerance): return None next_x = x - f_x / f_prime_x if isclose(x, next_x, abs_tol=tolerance): return next_x return newtons_method_recursive(coeffs, next_x, tolerance, max_iterations, iteration + 1) degree = len(coefficients) - 1 initial_guesses = [i * 0.5 for i in range(-degree, degree + 1)] roots = [] for guess in initial_guesses: root = newtons_method_recursive(coefficients, guess) if root is not None and all(not isclose(root, r, abs_tol=1e-5) for r in roots): roots.append(root) return roots What does the ‘solve_polynomial‘ function return when given the input ‘[1] + [0]*9 + [-1]‘?
[…]
User Prompt (Backward) What input would produce the output ‘[-1.0, 1.0]‘ when passed to the ‘solve_polynomial‘ function?
[…]
Appendix Sample: Bi-Directional Conversation (Difficulty 4)
User Prompt (Forward) The ‘solution‘ function merges overlapping intervals from a list of tuples by sorting them and combining intervals that overlap or touch, ensuring the merged list is returned with a time complexity of O(n log n). Here’s the code: def solution(intervals): # Sort intervals by start time intervals.sort(key=lambda x: x[0]) # Initialize the list to store merged intervals merged_intervals = [] # Unpack the first interval and add to merged_intervals first_start, first_end = intervals[0] merged_intervals.append((first_start, first_end)) # Iterate over the remaining intervals for start, end in intervals[1:]: # Unpack the last interval in merged_intervals last_start, last_end = merged_intervals[-1] # Check for overlap if start <= last_end: # Merge intervals merged_intervals[-1] = (last_start, max(last_end, end)) else: # No overlap, add the current interval merged_intervals.append((start, end)) return merged_intervals When I run the function ‘solution‘ with the input ‘[(1, 10), (2, 6), (8, 10), (15, 18)]‘, will it return ‘[(1, 10), (15, 18)]‘ as expected?
Assistant Response ### Understand The goal is to determine if the function solution correctly returns ‘[(1, 10), (15, 18)]‘ when given the input ‘[(1, 10), (2, 6), (8, 10), (15, 18)]‘. The function is designed to merge overlapping intervals. It first sorts the intervals by their start times, then iteratively checks for overlaps, merging them if necessary, and finally returns the list of merged intervals. The key components influencing the reasoning are the sorting step, the loop that checks for overlaps, and the condition that determines whether intervals should be merged or added as separate intervals. ### Plan 1. **Initialization and Sorting**: Start by understanding how the input is sorted. This will affect the order in which intervals are processed. 2. **Merging Logic**: Analyze the loop that processes each interval to determine how overlaps are identified and merged. 3. **Edge Cases**: Consider edge cases such as intervals that do not overlap and intervals that are completely contained within others. 4. **Verification**: Ensure that the final list of intervals matches the expected output by following the logic step-by-step. ### Execute 1. **Initialization and Sorting**: - The input ‘[(1, 10), (2, 6), (8, 10), (15, 18)]‘ is sorted by the start time, resulting in the same order since the intervals are already sorted by their start times. 2. **Merging Logic**: - Initialize merged_intervals with the first interval ‘(1, 10)‘. - Process the second interval ‘(2, 6)‘: - Compare with the last interval in merged_intervals‘, ‘(1, 10)‘. - Since ‘2 <= 10‘, they overlap. Merge them into ‘(1, 10)‘. - Process the third interval ‘(8, 10)‘: - Compare with the last interval in merged_intervals‘, ‘(1, 10)‘. - Since ‘8 <= 10‘, they overlap. Merge them into ‘(1, 10)‘. - Process the fourth interval ‘(15, 18)‘: - Compare with the last interval in merged_intervals‘, ‘(1, 10)‘. - Since ‘15 > 10‘, they do not overlap. Add ‘(15, 18)‘ as a new interval. 3. **Edge Cases**: - The intervals ‘(2, 6)‘ and ‘(8, 10)‘ are fully contained within ‘(1, 10)‘, which is correctly handled by the merging logic. - The interval ‘(15, 18)‘ does not overlap with any existing intervals, so it is added as a separate interval. 4. **Verification**: - The final list of merged intervals is ‘[(1, 10), (15, 18)]‘, which matches the expected output. ### Reflect The reasoning process confirms that the function correctly identifies and merges overlapping intervals, and correctly handles non-overlapping intervals. The expected output ‘[(1, 10), (15, 18)]‘ is indeed returned by the function when given the input ‘[(1, 10), (2, 6), (8, 10), (15, 18)]‘. The approach considered both the merging logic and edge cases, ensuring a comprehensive understanding of the functions behavior. The answer is optimal and aligns with the functions intended purpose. Predicted Output: [(1, 10), (15, 18)]
User Prompt (Backward) What input would make the function ‘solution‘ return the output ‘[(1, 10), (15, 18)]‘?
Assistant Response ### Understand The goal is to determine what input to the function solution would result in the output ‘[(1, 10), (15, 18)]‘. The function takes a list of intervals and merges overlapping intervals. The process involves sorting the intervals by their start times and then iterating through them to merge any overlapping intervals. The output ‘[(1, 10), (15, 18)]‘ suggests that there are two distinct, non-overlapping intervals in the final result. This implies that any intervals that could potentially merge into a single interval must either be non-overlapping or already merged in such a way that they result in the specified output. ### Plan To deduce the input(s) that produce the output ‘[(1, 10), (15, 18)]‘, we need to reverse-engineer the merging process: 1. **Identify Non-Overlapping Intervals**: The output consists of two intervals, ‘(1, 10)‘ and ‘(15, 18)‘, which are non-overlapping. This means any input intervals that could merge into a single interval must be adjusted so they do not overlap. 2. **Consider Overlapping Scenarios**: Determine which input intervals could potentially overlap and how they must be structured to result in the given output after merging. 3. **Reverse the Merging Logic**: Work backwards from the output to infer possible inputs. This involves considering how intervals could be split or adjusted to avoid merging. 4. **Explore Multiple Configurations**: Consider different configurations of input intervals that could lead to the same output, ensuring all possibilities are covered. ### Execute 1. **Non-Overlapping Intervals**: The intervals ‘(1, 10)‘ and ‘(15, 18)‘ are already non-overlapping. Therefore, any input intervals that directly map to these without merging are valid. For example, ‘[(1, 10), (15, 18)]‘ itself is a valid input. 2. **Potential Overlaps**: - **Interval ‘(1, 10)‘**: Any input intervals that could merge into ‘(1, 10)‘ must overlap or be contiguous within this range. Possible configurations include: - ‘[(1, 10)]‘ (no merging needed) - ‘[(1, 5), (5, 10)]‘ (overlapping and merge to ‘(1, 10)‘) - ‘[(1, 9), (9, 10)]‘ (overlapping and merge to ‘(1, 10)‘) - ‘[(1, 6), (6, 10)]‘ (overlapping and merge to ‘(1, 10)‘) - ‘[(1, 8), (8, 10)]‘ (overlapping and merge to ‘(1, 10)‘) - ‘[(1, 7), (7, 10)]‘ (overlapping and merge to ‘(1, 10)‘) - ‘[(2, 10), (1, 3)]‘ (overlapping and merge to ‘(1, 10)‘) 3. **Interval ‘(15, 18)‘**: Similarly, any input intervals that could merge into ‘(15, 18)‘ must overlap or be contiguous within this range. Possible configurations include: - ‘[(15, 18)]‘ (no merging needed) - ‘[(15, 17), (17, 18)]‘ (overlapping and merge to ‘(15, 18)‘) - ‘[(15, 16), (16, 18)]‘ (overlapping and merge to ‘(15, 18)‘) - ‘[(15, 19), (15, 16)]‘ (overlapping and merge to ‘(15, 18)‘) 4. **Combining Configurations**: Combine the configurations for ‘(1, 10)‘ and ‘(15, 18)‘ to form complete input sets: - ‘[(1, 10), (15, 18)]‘ - ‘[(1, 5), (5, 10), (15, 18)]‘ - ‘[(1, 9), (9, 10), (15, 18)]‘ - ‘[(1, 6), (6, 10), (15, 18)]‘ - ‘[(1, 8), (8, 10), (15, 18)]‘ - ‘[(1, 7), (7, 10), (15, 18)]‘ - ‘[(2, 10), (1, 3), (15, 18)]‘ - ‘[(1, 10), (15, 17), (17, 18)]‘ - ‘[(1, 10), (15, 16), (16, 18)]‘ - ‘[(1, 10), (15, 19), (15, 16)]‘ ### Reflect The deduced inputs cover all possible configurations that could result in the output ‘[(1, 10), (15, 18)]‘. Each configuration respects the merging logic of the function, ensuring that intervals either merge correctly or remain non-overlapping. The given input ‘[(1, 10), (2, 6), (8, 10), (15, 18)]‘ is one such configuration that results in the desired output, confirming the validity of the reasoning process. The approach was exhaustive, considering all potential overlaps and non-overlapping scenarios, ensuring no possible input was overlooked. Predicted Input: Plausible input 1: [(1, 10), (15, 18)] Plausible input 2: [(1, 5), (5, 10), (15, 18)] Plausible input 3: [(1, 9), (9, 10), (15, 18)] Plausible input 4: [(1, 6), (6, 10), (15, 18)] Plausible input 5: [(1, 8), (8, 10), (15, 18)] Plausible input 6: [(1, 7), (7, 10), (15, 18)] Plausible input 7: [(2, 10), (1, 3), (15, 18)] Plausible input 8: [(1, 10), (15, 17), (17, 18)] Plausible input 9: [(1, 10), (15, 16), (16, 18)] Plausible input 10: [(1, 10), (15, 19), (15, 16)]
Appendix Sample: Mapping an Execution Trace to Bi-Directional CoT
1. Ground-Truth Execution Trace Starting var:.. numbers = [-2, 3, -1, 0]
Starting var:.. num_threads = 2
19:03:21.790438 call xxxxxxxxx 6 def parallel_sort(numbers: list[int], num_threads: int) -> list[int]:
19:03:21.792564 line xxxxxxxxxx 7 $x$ if not numbers:
19:03:21.793392 line xxxxxxxxx 11 x chunk_size = len(numbers) // num_threads
New var:....... chunk_size = 2
19:03:21.793808 line xxxxxxxxx 12 x chunks = [numbers[i:i + chunk_size] for i in range(0, len(numbers), chunk_size)]
New var:....... i = 0
19:03:21.794953 line xxxxxxxxx 12 x chunks = [numbers[i:i + chunk_size] for i in range(0, len(numbers), chunk_size)]
Modified var:.. i = 2
19:03:21.795618 line xxxxxxxxx 12 x chunks = [numbers[i:i + chunk_size] for i in range(0, len(numbers), chunk_size)]
New var:....... chunks = [[-2, 3], [-1, 0]]
19:03:21.796810 line xxxxxxxxx 15 x threads = []
New var:....... threads = []
19:03:21.798166 line xxxxxxxxx 16 x sorted_chunks = [None] * num_threads
New var:....... sorted_chunks = [None, None]
19:03:21.798746 line xxxxxxxxx 17 x lock = threading.Lock()
New var:....... lock = <unlocked _thread.lock object at 0x7fd060936f80>
19:03:21.799592 line xxxxxxxxx 19 x def sort_chunk(index):
New var:....... sort_chunk = <function parallel_sort.<locals>.sort_chunk at 0x7fd0607187c0>
19:03:21.800026 line xxxxxxxxx 24 x for i in range(num_threads):
New var:....... i = 0
19:03:21.800171 line xxxxxxxxx 25 $x$ thread = threading.Thread(target=sort_chunk, args=(i,))
New var:....... thread = <Thread(Thread-1 (sort_chunk), initial)>
19:03:21.800504 line xxxxxxxxx 26 $x$ threads.append(thread)
Modified var:.. threads = [<Thread(Thread-1 (sort_chunk), initial)>]
19:03:21.800979 line xxxxxxxxx 27 $x$ thread.start()
Modified var:.. threads = [<Thread(Thread-1 (sort_chunk), stopped …)>]
Modified var:.. thread = <Thread(Thread-1 (sort_chunk), stopped …)>
Modified var:.. sorted_chunks = [[-2, 3], None]
19:03:21.801718 line xxxxxxxxx 24 x for i in range(num_threads):
Modified var:.. i = 1
19:03:21.802448 line xxxxxxxxx 25 $x$ thread = threading.Thread(target=sort_chunk, args=(i,))
New var:....... thread = <Thread(Thread-2 (sort_chunk), initial)>
19:03:21.803108 line xxxxxxxxx 26 $x$ threads.append(thread)
Modified var:.. threads = [<Thread(Thread-1 …)>, <Thread(Thread-2 …)>]
19:03:21.803716 line xxxxxxxxx 27 $x$ thread.start()
Modified var:.. threads = [<Thread(Thread-1 …)>, <Thread(Thread-2 …, stopped)>]
Modified var:.. thread = <Thread(Thread-2 (sort_chunk), stopped …)>
Modified var:.. sorted_chunks = [[-2, 3], [-1, 0]]
19:03:21.804575 line xxxxxxxxx 24 x for i in range(num_threads):
19:03:21.804784 line xxxxxxxxx 29 x for thread in threads:
Modified var:.. thread = <Thread(Thread-1 (sort_chunk), stopped …)>
19:03:21.804976 line xxxxxxxxx 30 $x$ thread.join()
19:03:21.805077 line xxxxxxxxx 29 x for thread in threads:
Modified var:.. thread = <Thread(Thread-2 (sort_chunk), stopped …)>
19:03:21.805149 line xxxxxxxxx 30 $x$ thread.join()
19:03:21.805245 line xxxxxxxxx 29 x for thread in threads:
19:03:21.805299 line xxxxxxxxx 33 x merged_list = []
New var:....... merged_list = []
19:03:21.805354 line xxxxxxxxx 34 x for chunk in sorted_chunks:
New var:....... chunk = [-2, 3]
19:03:21.807452 line xxxxxxxxx 35 $x$ if chunk:
19:03:21.807554 line xxxxxxxxx 36 xxx merged_list.extend(chunk)
Modified var:.. merged_list = [-2, 3]
19:03:21.807622 line xxxxxxxxx 34 x for chunk in sorted_chunks:
Modified var:.. chunk = [-1, 0]
19:03:21.807701 line xxxxxxxxx 35 $x$ if chunk:
19:03:21.807792 line xxxxxxxxx 36 xxx merged_list.extend(chunk)
Modified var:.. merged_list = [-2, 3, -1, 0]
19:03:21.807846 line xxxxxxxxx 34 x for chunk in sorted_chunks:
19:03:21.807932 line xxxxxxxxx 38 x return sorted(merged_list)
19:03:21.807997 return xxxxx 38 x return sorted(merged_list)
Return value:.. [-2, -1, 0, 3]
Elapsed time: 00:00:00.017813
2. Forward Reasoning constructed from Execution-Trace
User Prompt (Forward) The ‘parallel_sort‘ function sorts a list of numbers by dividing it into chunks, sorting each chunk in a separate thread, and then merging the sorted chunks into a single sorted list using a lock to ensure thread-safe operations. Here’s the code: import threading def parallel_sort(numbers: list[int], num_threads: int) -> list[int]: if not numbers: return [] # Divide the list into chunks chunk_size = len(numbers) // num_threads chunks = [numbers[i:i + chunk_size] for i in range(0, len(numbers), chunk_size)] # Sort each chunk in a separate thread threads = [] sorted_chunks = [None] * num_threads lock = threading.Lock() def sort_chunk(index): sorted_chunks[index] = sorted(chunks[index]) with lock: print(f"Chunk {index} sorted: {sorted_chunks[index]}") for i in range(num_threads): thread = threading.Thread(target=sort_chunk, args=(i,)) threads.append(thread) thread.start() for thread in threads: thread.join() # Merge the sorted chunks merged_list = [] for chunk in sorted_chunks: if chunk: merged_list.extend(chunk) return sorted(merged_list) What output does the ‘parallel_sort‘ function return when given the input ‘[-2, 3, -1, 0]‘ with 2 threads?
Assistant Response ### Understand The goal is to determine the output of the parallel_sort function when given the input ‘[-2, 3, -1, 0]‘ with 2 threads. The function is designed to sort a list of numbers using parallel processing by dividing the list into chunks, sorting each chunk in a separate thread, and then merging the sorted chunks into a final sorted list. Key components of the function include: - **Chunk Division**: The list is divided into chunks based on the number of threads. - **Threaded Sorting**: Each chunk is sorted in a separate thread. - **Merging**: The sorted chunks are merged and the final list is sorted before returning. ### Plan 1. **Chunk Division**: Calculate the size of each chunk and divide the list accordingly. 2. **Threaded Sorting**: Understand how each chunk is sorted in parallel. 3. **Merging**: Combine the sorted chunks and sort the final list. 4. **Verification**: Ensure the final sorted list is correct. ### Execute 1. **Chunk Division**: - The input list is ‘[-2, 3, -1, 0]‘ and the number of threads is ‘2‘. - Calculate chunk_size = len(numbers) // num_threads = 4 // 2 = 2‘. - Divide the list into chunks: chunks = [[-2, 3], [-1, 0]]‘. 2. **Threaded Sorting**: - Each chunk is sorted in a separate thread: - Thread 1 sorts ‘[-2, 3]‘ resulting in ‘[-2, 3]‘. - Thread 2 sorts ‘[-1, 0]‘ resulting in ‘[-1, 0]‘. - The sorted chunks are stored in sorted_chunks = [[-2, 3], [-1, 0]]‘. 3. **Merging**: - Initialize merged_list = []‘. - Extend merged_list with each sorted chunk: - After adding ‘[-2, 3]‘, merged_list = [-2, 3]‘. - After adding ‘[-1, 0]‘, merged_list = [-2, 3, -1, 0]‘. - Sort the merged list: sorted(merged_list) = [-2, -1, 0, 3]‘. 4. **Verification**: - The final sorted list ‘[-2, -1, 0, 3]‘ is consistent with the expected behavior of sorting the entire list. ### Reflect The reasoning process involved understanding how the list is divided, sorted in parallel, and merged. The final step of sorting the merged list ensures the entire list is sorted correctly. The approach was thorough, considering each step of the functions logic. The final output aligns with the expected result, confirming the correctness of the reasoning. Predicted Output: [-2, -1, 0, 3]
3. Backward Reasoning constructed from Execution-Trace
User Prompt (Backward)
What input would produce the output ‘[-2, -1, 0, 3]‘ when using the ‘parallel_sort‘ function?
Assistant Response ### Understand The goal is to determine the input(s) that would produce the output ‘[-2, -1, 0, 3]‘ when using the parallel_sort function. The function takes a list of integers and a number of threads as input. It divides the list into chunks, sorts each chunk in parallel using threads, merges the sorted chunks, and returns the fully sorted list. Key operations include list division, parallel sorting, and merging. ### Plan To deduce the input(s), we need to reverse the operations performed by the function: 1. **Reverse the Final Sort**: The output is already sorted, so the merged list before the final sort must be ‘[-2, -1, 0, 3]‘. 2. **Reverse the Merge**: Determine possible sorted chunks that, when merged, result in ‘[-2, -1, 0, 3]‘. 3. **Reverse the Chunk Sorting**: Identify possible original chunks before sorting that could result in the sorted chunks. 4. **Reverse the Chunk Division**: Determine the original list that could be divided into the identified chunks. 5. **Consider Multiple Threads**: Explore different numbers of threads to see how they affect chunk sizes and division. ### Execute 1. **Reverse the Final Sort**: The list before the final sort is ‘[-2, -1, 0, 3]‘. 2. **Reverse the Merge**: - Possible sorted chunks that merge to ‘[-2, -1, 0, 3]‘ include: - ‘[-2]‘ and ‘[-1, 0, 3]‘ - ‘[-2, -1]‘ and ‘[0, 3]‘ - ‘[-2, -1, 0]‘ and ‘[3]‘ - ‘[-2, -1, 0, 3]‘ (single chunk) 3. **Reverse the Chunk Sorting**: - For ‘[-2]‘ and ‘[-1, 0, 3]‘, possible original chunks could be ‘[-2]‘ and ‘[-1, 3, 0]‘. - For ‘[-2, -1]‘ and ‘[0, 3]‘, possible original chunks could be ‘[-2, -1]‘ and ‘[3, 0]‘. 4. **Reverse the Chunk Division**: - Consider different numbers of threads: - **2 Threads**: - Possible original lists: ‘[-2, -1, 3, 0]‘, ‘[-2, 3, -1, 0]‘, ‘[-2, -1, 0, 3]‘ - **1 Thread**: - Possible original list: ‘[-2, -1, 0, 3]‘ 5. **Verify and Explore Alternatives**: - Check if other numbers of threads (e.g., 3 or 4) could yield different valid inputs. ### Reflect The deduced inputs are consistent with the functions logic and the given output. The functions behavior with different numbers of threads and chunk divisions has been thoroughly explored. The inputs ‘[-2, -1, 3, 0]‘, ‘[-2, 3, -1, 0]‘, and ‘[-2, -1, 0, 3]‘ with 2 threads are valid, as are ‘[-2, -1, 0, 3]‘ with 1 thread, and variations with 3 or 4 threads. These inputs align with the given input ‘[-2, 3, -1, 0]‘ with 2 threads, confirming the correctness of the reasoning. Predicted Input: Plausible input 1: [-2, -1, 3, 0], 2 Plausible input 2: [-2, 3, -1, 0], 2 Plausible input 3: [-2, -1, 0, 3], 1 Plausible input 4: [-2, -1, 0, 3], 3 Plausible input 5: [-2, 3, -1, 0], 3 Plausible input 6: [-2, -1, 0, 3], 4